### v1

# Parabolic bundles and the intersection cohomology of moduli spaces of vector bundles on curves

Abstract.
The study of the intersection cohomology of moduli spaces of semistable bundles was initiated by Frances Kirwan in the 1980’s. In this paper, we give a complete geometric proof of a recursive formula, which reduces the calculation of the intersection Betti numbers of the moduli spaces of semistable bundles on Riemann surfaces in degree-0 and arbitrary rank to the known formulas of the Betti numbers of the smooth, degree-1 moduli spaces. Our formula was motivated by the work of Mozgovoy and Reineke in 2015, and appears as a consequence of the Decomposition Theorem applied to the forgetful map from a parabolic moduli space. Our methods are direct, geometric and use a minimal amount of background.
1. Introduction: the intersection cohomology of moduli spaces
Let $C$ be a smooth complex projective curve of genus $g\geq 2$, and denote by $\mathcal{M}_{d}(r)$ the moduli space of semistable degree-$d$, rank-$r$ vector bundles on $C$. When $r$ and $d$ are coprime, $\mathcal{M}_{d}(r)$ is a smooth projective variety whose cohomology has been computed using a number of techniques, starting in the 1970’s [16, 10, 1, 12, 3]. However, when $\mathrm{gcd}(d,r)>1$, the moduli space $\mathcal{M}_{d}(r)$ is singular, and, in this case, a natural enumerative invariant is intersection cohomology, introduced by Goresky and MacPherson in the late 1970’s [15]. The project to calculate this ”new” cohomology for the moduli spaces $\mathcal{M}_{d}(r)$ of semistable bundles on Riemann surfaces was initiated by Frances Kirwan not long afterwards [18, 20, 19]. One of the most interesting cases is $\mathcal{M}_{0}(r)$ for $r>1$, which is diffeomorphic to the character variety of unitary representations of the fundamental group of $C$ [27]. Over the years, a lot of progress was made, but mostly in the rank-2 case, until in 2015, Mozgovoy and Reineke wrote down a general formula for the intersection Betti numbers of $\mathcal{M}_{d}(r)$ in all ranks and degrees [25]. They used an argument from Donaldson–Thomas theory, which had a gap, however.  [^1] The gap has apparently been fixed recently.
In the present paper, we calculate the intersection Betti numbers of $\mathcal{M}_{0}(r)$ in all ranks using the Decomposition Theorem applied to a particular projection map onto $\mathcal{M}_{0}(r)$, and drawing on some ideas from [32]. We paint a detailed geometric picture of the geometry of parabolic structures on bundles, giving access to Hodge structures, intersection numbers, and other data of the theory. Along the way, we present a fresh look on the Decomposition Theorem itself.
In this introductory chapter, we briefly describe the basic properties of intersection cohomology that we will need, and then give a brisk introduction to the moduli spaces of semistable vector bundles, which is just sufficient to be able to present our main result, Theorem 1.4. This is followed by a description of our strategy, based on a version of the Decomposition Theorem adapted to our needs. For the reader with less patience, the rank-2 case is presented in detail in the following chapter §2.
Acknowledgments The authors wish to thank Ben Davison, Mark De Cataldo, Tamas Hausel, Anton Mellit, Luca Migliorini and Vivek Shende for useful discussions.
[…]
1.1. Intersection cohomology
We begin with a quick introduction to intersection cohomology (cf. [7, 15]). Some of what follows is true in greater generality, but to simplify our discussion, we always assume that our space $Y$ is a (possibly singular) complex quasiprojective algebraic variety.
The rational intersection cohomology groups of $Y$ form a graded vector space $IH^{\bullet}(Y)=\oplus_{i\in\mathbb{Z}}IH^{i}(Y)$ such that
  1. (i)
    $IH^{i}(Y)$ is a finite dimensional $\mathbb{Q}$- vector space and
    $IH^{i}(Y)=0\text{ for }i\not\in\{0,\ldots,2\dim Y\};$
  2. (ii)
    $IH^{\bullet}(Y)$ coincides with usual cohomology $H^{\bullet}(Y)$ when $Y$ has at worst finite quotient singularities.
  3. (iii)
    If $\widetilde{Y}$ is a normalization of $Y$, then $IH^{\bullet}(\widetilde{Y})\cong IH^{\bullet}(Y)$.
  4. (iv)
    (Hard Lefschetz) If $Y$ is projective and $\eta$ is an ample class on $Y$, then there is an isomorphism
[…]
  5. (v)
    (Poincaré-Verdier duality) there is a non-degenerate Poincaré pairing on $IH^{\bullet}(Y)$:
    $IH^{i}(Y)\otimes IH^{2\dim Y-i}(Y)\to\mathbb{Q}$
    for all $i\leq 2\dim Y$.
1.2. IC-sheaves
In what follows, $\mathbb{Q}_{Y}$ will stand for the constant sheaf on the variety $Y$. Just as usual cohomology may be described as the cohomology of the constant sheaf: $H^{\bullet}(Y)=\mathbb{H}^{\bullet}(Y,\mathbb{Q}_{Y})$, there is a (constructible) complex of sheaves $IC_{Y}$ on $Y$, called the IC-sheaf of $Y$, whose hypercohomology gives the intersection cohomology groups:
$IH^{i}(Y)\overset{\mathrm{def}}{=}\mathbb{H}^{i}(Y,IC_{Y}).$
Convention: In this paper, we set the cohomology sheaves of the complex $IC_{Y}$ to be graded from 0 to $2\dim Y$, so that if $Y$ is nonsingular we have $IC_{Y}=\mathbb{Q}_{Y}$.
As in the case of ordinary cohomology, intersection cohomology admits coefficients in local systems: assume that $Y$ is (Whitney) stratified by smooth quasi-projective varieties:
$Y=\sqcup_{\alpha}Y_{\alpha};$
given a local system $\mathcal{L}$ on a stratum $Y_{\alpha}$, there is a general construction
$\mathcal{L}\rightsquigarrow IC(\overline{Y}_{\alpha},\mathcal{L}),$
of a constructible complex $IC(\overline{Y}_{\alpha},\mathcal{L})$ on $Y$ called the intermediate extension of $\mathcal{L}$. This complex is supported on the closure $\overline{Y}_{\alpha}$, restricts to $\mathcal{L}$: $IC(\overline{Y}_{\alpha},\mathcal{L})|_{Y_{\alpha}}=\mathcal{L}$, and satisfies some additional support conditions (see [15]). The corresponding twisted intersection cohomology groups will be denoted as $IH^{i}(\overline{Y}_{\alpha},\mathcal{L})\overset{\mathrm{def}}{=}\mathbb{H}^{i}(Y,IC(\overline{Y}_{\alpha},\mathcal{L})).$ In particular, $IC_{Y}=IC(Y,\mathbb{Q}_{Y_{0}})$, where $Y_{0}$ is the top-dimensional stratum of smooth points of $Y$.
While in general the intersection cohomology of a variety $Y$ is not easy to determine, we can describe it explicitly when $Y$ is an affine cone over a projective variety (see [23, 11]).
Proposition 1.1 (Intersection cohomology of an affine cone).
Let $Z$ be a complex projective variety of dimension $n-1$, endowed with an ample line bundle $\eta$. Let $Y$ be the affine cone over $Z$ given by contracting the zero-section in the total space of $\eta$. Then
$IH^{d}(Y)\cong\begin{cases}IH_{\mathrm{prim}}^{d}(Z)&\mbox{ if $d<n$}\\ 0&\mbox{ otherwise,}\end{cases}$
where the primitive cohomology is given by (see the Hard Lefschetz property above)
$IH^{d}_{\mathrm{prim}}(Z)=\operatorname{ker}\left\{\cup c_{1}(\eta)^{n-d+1}:IH^{d}(Z)\rightarrow IH^{2n-d+2}(Z)\right\}.$
Conventions regarding local systems: Throughout the paper, given a morphism $f:X\rightarrow Y$ of algebraic varieties and an open set $U\subseteq Y$ over which $f$ is a fibration, we will assume that the local systems $R^{j}f_{*}\mathbb{Q}_{f^{-1}(U)}$ are placed in degree $j$ when considered as constructible complexes on $U$. Similarly, we will write $\mathcal{L}^{j}$ for a local system shifted in degree $j$ and, with some abuse of terminology, still refer to a direct sum $\mathcal{L}=\bigoplus_{j}\mathcal{L}^{j}$ as a local system.
In this case, we set
$IC(Y,\mathcal{L}):=\bigoplus_{j}IC(Y,\mathcal{L}^{j}).$
The goal of this article is to calculate the intersection cohomology of certain moduli spaces. The methods we employ are indirect, and we will explain them in a later part of this introduction. We now describe the topological spaces in question.
1.3. Semistable bundles
Let $C$ be a smooth complex projective curve of genus $g\geq 2$. For a vector bundle $V$ on $C$, define the slope of $V$ as
$\mu(V)=\frac{\mathrm{deg}(V)}{\mathrm{rk}(V)}.$
  • We say that $V$ is semistable if for all proper subbundles $V^{\prime}\subset V$ we have
    (1) $\mu(V^{\prime})\leq\mu(V),\text{ and }$
    (2) $V\text{ is {stable} if }V^{\prime}\subset V\Rightarrow\mu(V^{\prime})<\mu(V).$
    Obviously, if the degree and the rank are coprime, then semistability implies stability.
  • We say that $V$ is polystable if it is isomorphic to a direct sum of stable vector bundles with equal slopes. Every polystable bundle is semistable.
  • Every semistable bundle $V$ on $C$ admits a Jordan-Hölder filtration with stable subquotients, i.e. there exists a sequence of vector bundles
    $0=W_{0}\subseteq W_{1}\subseteq\ldots\subseteq W_{r}=V,$
    where $V_{i}=W_{i}/W_{i-1}$ are stable bundles with $\mu(V_{i})=\mu(V)$. Although the Jordan-Hölder filtration is not unique, the quotients up to permutation are, hence we can associate to a semistable bundle $V$ a unique polystable bundle:
    $V\mapsto\mathrm{Gr}(V)\overset{\mathrm{def}}{=}\bigoplus_{i}V_{i}.$
  • We say that vector bundles $V$ and $V^{\prime}$ are s-equivalent, and write $V\sim_{s}V^{\prime}$ if
    $\mathrm{Gr}(V)\cong\mathrm{Gr}(V^{\prime}).$
Using Geometric Invariant Theory, Mumford constructed a coarse moduli space $\mathcal{M}_{d}(r)$ parametrizing s-equivalence classes of rank $r$ and degree $d$ vector bundles on $C$ (see [26]).
Theorem 1.2.
Let $C$ be a smooth complex projective curve of genus $g\geq 2$. Then there is a projective algebraic variety $\mathcal{M}_{d}(r)$ of dimension $(g-1)r^{2}+1$, which serves as a coarse moduli space parametrizing rank-$r$, degree-$d$ semistable bundles up to $s$-equivalence.
  • As a set $\mathcal{M}_{d}(r)$ is in one-to-one correspondence with the set of (isomoprphism classes) of polystable bundles;
  • the subset $\mathcal{M}_{d}(r)^{\mathrm{st}}$ of $\mathcal{M}_{d}(r)$ parametrizing stable bundles up to isomorphism is a smooth, Zariski open subvariety; in particular, $\mathcal{M}_{d}(r)$ is smooth if $\mathrm{gcd}(r,d)=1$. In this case, there exists a universal bundle $U_{d}(r)$ over $\mathcal{M}_{d}(r)\times C$, satisfying the standard tautological properties.
The Poincaré polynomials of $\mathcal{M}_{d}(r)$ for $\mathrm{gcd}(r,d)=1$ have been computed in several ways, see [16, 10, 1, 12]. In this article, we express the intersection Betti numbers of the singular moduli spaces $\mathcal{M}_{0}(r)$, for all $r\geq 1$, via the Poincaré polynomials for $\mathcal{M}_{1}(r)$. Our formula uses an algebraic device called the plethystic exponential, which we will now describe.
1.4. Graded vector spaces and the plethystic exponential
1.4.1. Notation
  • For a graded vector space $\mathcal{H}^{\bullet}=\oplus_{i}\mathcal{H}^{i}$, with finite-dimensional components, introduce the notation
    $\mathrm{Hilb}_{t}(\mathcal{H}^{\bullet})\overset{\mathrm{def}}{=}\sum_{i}t^{i}\dim\mathcal{H}^{i}$
    for its Hilbert series. In particular, for an oriented manifold $M$,
    $P_{t}(M)=\mathrm{Hilb}_{t}(H^{\bullet}(M))$
    is its Poincaré polynomial, while if $M$ is an algebraic variety, then we introduce
    $IP_{t}(M)=\mathrm{Hilb}_{t}(IH^{\bullet}(M)).$
  • Similarly, we set
    $\mathrm{Hilb}_{t,q}(\mathcal{H}^{\bullet})\overset{\mathrm{def}}{=}\sum_{i,j}t^{i}q^{j}\dim\mathcal{H}^{ij}$
    for a bigraded vector space $\mathcal{H}^{\bullet}=\oplus_{ij}\mathcal{H}^{ij}$.
1.4.2. The plethystic exponential
Consider $\mathbb{Z}^{d}\subset\mathbb{R}^{d}$ as the lattice of possible exponents $\eta$ of monomials $m^{\eta}$, and let $U$ be an open halfspace in $\mathbb{R}^{d}$ containing the origin in its closure. Given a Laurent polynomial (or formal power series) with integer coefficients whose exponent set is supported in $U$
$Q=\sum_{\eta\in U\cap\mathbb{Z}^{d}}a_{\eta}m^{\eta},\quad\text{where }a_{\eta}\in\mathbb{Z}\;\forall\eta\in U\cap\mathbb{Z}^{d},$
we can form the fraction
(3) $\mathbb{E}\mathrm{xp}[Q]=\prod_{\eta\in U\cap\mathbb{Z}^{d}}(1-m^{\eta})^{-a_{\eta}},$
which may be expanded into a Laurent power series with support in $\{0\}\cup(U\cap\mathbb{Z}^{d})$. This operation is known as the plethystic exponential, which is a group homomorphism from the additive group of Laurent series with support in $U$ to the multiplicative group of functions of the form $1+$ such a Laurent series.
In general, the plethystic exponential is a rather intractable object, but with some parity restrictions, it has a simple interpretation. Introduce the free $\mathbb{Z}/2$-commutative ring
(4) $\mathbb{Q}_{[t]}[\mathcal{H}^{\bullet}]=\mathrm{Sym}^{\bullet}\left(\mathcal{H}_{[t]}^{\mathrm{even}}\right)\bigotimes\wedge^{\bullet}\left(\mathcal{H}_{[t]}^{\mathrm{odd}}\right),$
where
$\mathcal{H}_{[t]}^{\mathrm{even}}=\bigoplus_{i\,\mathrm{even}}\mathcal{H}^{ij},\quad\mathcal{H}_{[t]}^{\mathrm{odd}}=\bigoplus_{i\,\mathrm{odd}}\mathcal{H}^{ij}.$
Lemma 1.3.
Let $\mathcal{H}^{\bullet}$ be a bigraded vector space with finite-dimensional components as above. Then
$\mathrm{Hilb}_{-t,q}(\mathbb{Q}_{[t]}[\mathcal{H}^{\bullet}])=\mathbb{E}\mathrm{xp}[\mathrm{Hilb}_{-t,q}(\mathcal{H}^{\bullet})].$
The proof of this statement hinges on the fact that, when expanding the fraction in (3) for $\mathrm{Hilb}_{-t,q}(\mathcal{H}^{\bullet})$, the contributions to a particular bigraded piece all have the same signs, and thus there are no cancellations between positive and negative terms.
1.5. The main result
Our main result is the following formula.
Theorem 1.4.
The Poincaré polynomials of the smooth spaces $\mathcal{M}_{1}(r)$ and the intersection Poincaré polynomials of the singular spaces $\mathcal{M}_{0}(r)$, $r=1,2,\dots$ are related as follows:
(5) $1+\sum_{r=1}^{\infty}P_{-t}(\mathbb{P}^{r-1})\cdot P_{-t}(\mathcal{M}_{1}(r))\cdot(-t)^{(1-g)r^{2}}q^{r}=\\ =\mathbb{E}\mathrm{xp}\left[\sum_{r=1}^{\infty}P_{-t}(\mathbb{P}^{r-1})\cdot IP_{-t}(\mathcal{M}_{0}(r))\cdot(-t)^{(1-g)r^{2}}q^{r}\right].$
Remark 1.5.
Clearly, formula (5) provides us with a recursion for $IP_{t}(\mathcal{M}_{0}(r))$, if $P_{t}(\mathcal{M}_{1}(r))$ is known.
Remark 1.6.
Our main results, and the methods used have several implications.
  • Formula (5), may be refined to give the intersection Hodge numbers of $\mathcal{M}_{0}(r)$ (see (85)).
  • The proof of our formula allows to embed $IH^{\bullet}(\mathcal{M}_{0}(r))$ in $H^{\bullet}(\mathcal{P}_{1}(r))$ a canonical subspace, which gives us direct access to the intersection numbers of $\mathcal{M}_{0}(r)$.
1.6. The Decomposition Theorem
As mentioned in §1.2, direct calculations of intersection cohomologies are difficult. A more effective strategy is the use of the Decomposition Theorem [2] applied to proper maps $f:X\to Y$, where $X$ is smooth. Traditionally, the Decomposition Theorem is formulated in terms of the semisimplicity of the abelian category of perverse sheaves, and the fact that the simple objects in this category are the intermediate extensions of semisimple local systems.
We will use these results, but instead of investigating the perverse filtration, we will focus on refined methods of calculating the relevant local systems. The Theorem below is a somewhat naive version of the Decomposition Theorem, adapted to our needs. For a more detailed account on the Decomposition Theorem and, broadly, perverse sheaves we refer to De Cataldo and Migliorini [7], who provided a proof via classical Hodge theory in their seminal paper [6], and to the Bourbaki exposé by Williamson [33]
Theorem 1.7 (Decomposition theorem).
Let $f:X\to Y$ be a proper map from a smooth variety $X$ to a possibly singular $Y$, and assume that $Y=\sqcup_{\alpha}Y_{\alpha}$ is a stratification for $f$ in the sense that $f^{-1}(Y_{\alpha})$ is a fibration over $Y_{\alpha}$ for each $\alpha$. Let $p\in Y_{\beta}$, $U\subset Y$ a small open neighborhood of $p$, and consider the cohomological restriction map
(6) $j_{p}^{*}:H^{\bullet}(f^{-1}(U))\to H^{\bullet}(f^{-1}(U\setminus Y_{\beta})),\quad\text{where }j_{p}:f^{-1}(U\setminus Y_{\beta})\hookrightarrow f^{-1}(U)$
is the inclusion. Set $L_{p}:=\ker j_{p}^{*}$, and introduce the local system
(7) $\mathcal{L}_{\beta}\longrightarrow Y_{\beta}\text{ with fiber }L_{p},\,p\in Y_{\beta}.$
Then
  1. 0.
    $L_{p}$ is a graded ideal in $H^{\bullet}(f^{-1}(U))\cong H^{\bullet}(f^{-1}(p))$, whose graded pieces are in the degree range $[\dim X-\dim Y_{\beta}-r(f),\dim X-\dim Y_{\beta}+r(f)]$, where
    $r(f)=\dim X\times_{Y}X-\dim X$ is the defect of semismallness of $f$;
  2. I.
    in the derived category of bounded constructible complexes, there is a (non-canonical) direct sum decomposition
    (1.7.A) $Rf_{*}\mathbb{Q}_{X}\cong\bigoplus_{\alpha}IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha}).$
  3. II.
    In particular, passing to the cohomology sheaves of the complexes on both sides of this decomposition, and comparing the stalks of these sheaves, we have the isomorphism of vector spaces
    (1.7.B) $H^{\bullet}(f^{-1}(p))\cong L_{p}\oplus\bigoplus_{Y_{\beta}\in\overline{Y}_{\alpha}}\mathcal{H}^{\bullet}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha})\right)_{p}.$
  4. III.
    Finally, taking global cohomology over $Y$, we obtain the decomposition
    (1.7.C) $H^{\bullet}(X)\cong\bigoplus_{\alpha}IH^{\bullet}(\overline{{Y}}_{\alpha},\mathcal{L}_{\alpha}).$
The strata $Y_{\alpha}$ appearing in (1.7.A) for which $\mathcal{L}_{\alpha}$ does not vanish are called supports of $f$.
1.7. A strategy for calculating $IH(Y)$
We will now describe, in very general terms, how Theorem 1.7 allows one to calculate the intersection cohomology of $Y$.
  1. i.
    One needs to start with (1.7.B): first, one calculates the cohomologies of the fibers $f^{-1}(p)$ for all $p\in Y$, as well as the stalks of the IC sheaves appearing on the right hand side.
  2. ii.
    Armed with these quantities, one can set up a recursive calculation (the subtraction algorithm), which computes all the local system $\mathcal{L}_{\alpha}$ for each stratum $Y_{\alpha}$.
  3. iii.
    Now assume that $Y_{0}$ is simply connected; this implies that the local system $\mathcal{L}_{0}$ on $Y_{0}$ is trivial, and may be identified with the cohomology of the generic fiber $F_{0}$ of the map $f$. Then the contribution of $Y_{0}$ on the right hand side of (1.7.C) is
    (8) $IH^{\bullet}(Y,\mathcal{L}_{0})\cong\bigoplus_{i}IH^{\bullet}(Y,\mathbb{Q}_{Y}^{\oplus\dim H^{i}(F_{0})})\cong IH^{\bullet}(Y)\otimes H^{\bullet}(F_{0}),$
    and we need to subtract the contributions of the other strata from the Poincaré polynomial $P_{t}(X)$ in order to isolate (8). In fortunate situations, such as ours, these subtractions lead to another recursion, which, in our case, may be written in the form (5).
In order to describe our strategy precisely, we need to introduce some more notation and detail, so we postpone this until §3.6.
The content of the paper is as follows:
  • In §2, we present our results and the basics of our method for the case of $r=2$.
  • Next, in §3 we give an appropriate, natural stratification of our moduli spaces $\mathcal{M}_{0}(r)$, and construct the map $\pi:\mathcal{P}_{0}(r)\to\mathcal{M}_{0}(r)$ to which we apply the strategy described above. At the end of this Section in §3.6, we detail the implementation of the above strategy in our case.
  • In §4 we calculate the fiber cohomologies (cf. i. above).
  • In §5, 6,7 we perform the calculation of the local systems (cf. ii.).
  • Finally, in §8, we perform the recursion mentioned in iii. above.
2. The case of rank 2
While not all complications appear in the $r=2$ case, it is instructive to see our answer explicitly in this case. The final result (10) was first obtained by Kirwan [19].
2.1. The formula in rank 2
Comparing the coefficients of $q^{2}$ in (5), we obtain a formula of the form
$(1+t^{2})P_{-t}(\mathcal{M}_{1}(2))=(1+t^{2})IP_{-t}(\mathcal{M}_{0}(2))+t^{2(g-1)}\cdot[\text{contribution of }r=1].$
As we see from (3), this latter contribution depends crucially on the sign of $t$ on the right hand side of (5). This sign is $+1$ if $g$ is odd, and $-1$ if $g$ is even, and thus the contribution of $r=1$ is
$\begin{cases}\mathrm{Hilb}_{t}(\mathrm{Sym}^{2}H^{\mathrm{even}}\oplus\bigwedge^{2}H^{\mathrm{odd}}\oplus(H^{\mathrm{even}}\otimes H^{\mathrm{odd}}))\;\text{ if }g\text{ is odd,}\\ \mathrm{Hilb}_{t}(\bigwedge^{2}H^{\mathrm{even}}\oplus\mathrm{Sym}^{2}H^{\mathrm{odd}}\oplus(H^{\mathrm{even}}\otimes H^{\mathrm{odd}}))\;\text{ if }g\text{ is even.}\end{cases}$
As $\mathcal{M}_{0}(1)$ is the Jacobian variety of $C$, we have $P_{-t}(\mathcal{M}_{0}(1))=(1-t)^{2g}$, and, using basic character theory, we arrive at
(9) $(1+t^{2})P_{-t}(\mathcal{M}_{1}(2))=\\ (1+t^{2})IP_{-t}(\mathcal{M}_{0}(2))+\frac{1}{2}t^{2(g-1)}\left[(1-t)^{4g}+(-1)^{g-1}(1-t^{2})^{2g}\right],$
or, slightly rearranging the terms,
(10) $IP_{-t}(\mathcal{M}_{0}(2))=P_{-t}(\mathcal{M}_{1}(2))-\frac{t^{2(g-1)}}{2(1+t^{2})}\left[(1-t)^{4g}+(-1)^{g-1}(1-t^{2})^{2g}\right].$
2.2. The classification of semistable bundles in rank 2
The moduli space $\mathcal{M}_{0}(2)$ is stratified by 3 subvarieties [^2] The motivation for the notation will be given later, in §3. (cf. Theorem 1.2):
  1. (1)
    $M_{[2]}$, the smooth open stratum of stable bundles of $\dim=4g-3$,
  2. (2)
    $M_{[1,1]}^{\mathrm{ab}}$, the set $\{L_{1}\oplus L_{2}|\,L_{1}\not\cong L_{2}\in\mathcal{M}_{0}(1)\}$ of $\dim=2g$,
  3. (3)
    $M_{[1,1]}^{\mathrm{nab}}$, the locus $\{L\oplus L|\,L\in\mathcal{M}_{0}(1)\}$ of $\dim=g$.
The set of isomorphism classes of semistable bundles which are $s$-equivalent to a point in each stratum looks as follows:
  1. (1)
    just one element, the bundle $V$ itself;
  2. (2)
    the set of non-trivial extensions parameterized by the points of $\mathbb{P}\mathrm{Ext}^{1}(L_{1},L_{2})$ and $\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1})$, which are two copies of $\mathbb{P}^{g-2}$, and the single element, the polystable bundle $L_{1}\oplus L_{2}$ itself;
  3. (3)
    the set of non-trivial extensions parameterized by $\mathbb{P}\mathrm{Ext}^{1}(L,L)$, which is a copy of $\mathbb{P}^{g-1}$, and the polystable bundle $L\oplus L$ itself.
2.3. The fibers of the map $\pi$
We fix a point $p\in C$ and consider a smooth projective variety
$\mathcal{P}_{0}(2)=\{(V,F)|\,V-\text{semistable},\deg V=0,\mathrm{rk}V=2,F\subset V_{p}\text{ line},\\ [L\subsetneq V,\deg L=0\Longrightarrow L_{p}\neq F]\}/\cong$
of dimension $1+\dim\mathcal{M}_{0}(2)$ (cf. §3.3). Our plan is to analyze the Decomposition Theorem for the forgetful map $\pi:\mathcal{P}_{0}(2)\to\mathcal{M}_{0}(2)$. It is easy to check that the fibers of $\pi$ are as follows (see Figure 1):
  1. (1)
    if $V$ is stable, then $\pi^{-1}(V)$ may be identified with $\mathbb{P}(V_{p})\cong\mathbb{P}^{1}$.
  2. (2)
    $\pi^{-1}(L_{1}\oplus L_{2})$ when $L_{1}\not\cong L_{2}$, will consist of
    • a line bundle over $\mathbb{P}\mathrm{Ext}^{1}(L_{1},L_{2})$, corresponding to the extra data of $F\subset V_{p}\setminus(L_{2})_{p}$,
    • similarly, a line bundle over $\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1})$,
    • and a single point corresponding to $L_{1}\oplus L_{2}$.
    Altogether, this gives us two copies of $\mathbb{P}^{g-1}$ glued together along a single point. Note that the cohomology ring of this space is
    (11) $H^{\bullet}(\pi^{-1}(L_{1}\oplus L_{2}))\cong\mathbb{Q}[x_{1},x_{2}]/\langle x_{1}x_{2},x_{1}^{g},x_{2}^{g}\rangle.$
  3. (3)
    $\pi^{-1}(L\oplus L)\cong\mathbb{P}\mathrm{Ext}^{1}(L,L)$, which is a copy of $\mathbb{P}^{g-1}$.
[…]
Figure 1. The fibers of $\pi:\mathcal{P}_{0}(2)\to\mathcal{M}_{0}(2)$.
In section §4, we will give a more systematic and transparent description of the fibers in any rank.
2.4. The intersection cohomology of the normal cone for $M^{\mathrm{ab}}_{[1,1]}$.
Let $L_{1}\oplus L_{2}\in M^{\mathrm{ab}}_{[1,1]}$. The normal slice at this point is the affine GIT quotient
$\mathcal{C}_{L_{1}\oplus L_{2}}\cong\mathrm{Ext}^{1}(L_{1},L_{2})\oplus\mathrm{Ext}^{1}(L_{2},L_{1})\sslash\mathbb{C}^{*}\times\mathbb{C}^{*},$
where $\mathbb{C}^{*}\times\mathbb{C}^{*}$ is the automorphism group of $L_{1}\oplus L_{2}$. This space is a cone over the smooth variety
$\hat{\mathcal{C}}_{L_{1}\oplus L_{2}}\cong\mathbb{P}\mathrm{Ext}^{1}(L_{1},L_{2})\times\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1}),$
obtained by contracting the zero-section of the line bundle $\mathcal{O}(1)\boxtimes\mathcal{O}(1)$ to a point. Note that $\hat{\mathcal{C}}_{L_{1}\oplus L_{2}}$ has a natural action of the permutation group $\Sigma_{2}$, exchanging $L_{1}$ and $L_{2}$.
The stalk of the cohomology sheaves of $IC(\overline{M}_{[2]},\mathbb{Q})$ at $L_{1}\oplus L_{2}$ is the intersection cohomology of $\mathcal{C}_{L_{1}\oplus L_{2}}$, whose calculation is rather simple in this case: by Proposition 1.1 it is the primitive cohomology $H_{\mathrm{prim}}^{\bullet}(\hat{\mathcal{C}}_{L_{1}\oplus L_{2}})$ of $\hat{\mathcal{C}}_{L_{1}\oplus L_{2}}$ with respect to multiplication by $c_{1}(\mathcal{O}(1)\boxtimes\mathcal{O}(1))$. Explicitly, let $x_{1}=c_{1}(\mathcal{O}(1)\to\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1}))$ and $x_{2}=c_{1}(\mathcal{O}(1)\to\mathbb{P}\mathrm{Ext}^{1}(L_{1},L_{2}))$; then
$H^{\bullet}(\hat{\mathcal{C}}_{L_{1}\oplus L_{2}})\cong\mathbb{Q}[x_{1},x_{2}]/\langle x_{1}^{g-1},x_{2}^{g-1}\rangle\text{ and }H_{\mathrm{prim}}^{\bullet}(\hat{\mathcal{C}}_{L_{1}\oplus L_{2}})\cong H^{\bullet}(\hat{\mathcal{C}}_{L_{1}\oplus L_{2}})/\left\langle x_{1}+x_{2}\right\rangle.$
We thus obtain that $IH^{\bullet}(\mathcal{C}_{L_{1}\oplus L_{2}})$ has a ring structure in this case:
(12) $H_{\mathrm{prim}}^{\bullet}(\hat{\mathcal{C}}_{L_{1}\oplus L_{2}})\cong\mathbb{Q}[x_{1}-x_{2}]/\langle(x_{1}-x_{2})^{g-1}\rangle$
endowed with the action of the permutation group $\Sigma_{2}$ exchanging $x_{1}$ and $x_{2}$.
2.5. The subtraction algorithm
The equalities (1.7.B) and (1.7.C) of the Decomposition Theorem give rise to a subtraction algorithm, which we now apply to our map $\pi:\mathcal{P}_{0}(2)\to\mathcal{M}_{0}(2)$. We will then obtain (9) by taking the Hilbert functions on both sides of (1.7.C).
Using the Hecke correspondence (Proposition 3.11), we can conclude that there is a $\mathbb{P}^{1}$-fibration $\pi_{1}:\mathcal{P}_{0}(2)\to\mathcal{M}_{1}(2)$, which implies that
$P_{-t}(\mathcal{P}_{0}(2))=P_{-t}(\mathcal{M}_{1}(2)\times\mathbb{P}^{1}).$
This is exactly the expression which appears on the left hand side of (9). The right hand side of (9) is a priori a sum of 3 terms corresponding to the 3 strata listed above. To calculate them, we need to
  • first, identify the local systems $\mathcal{L}_{[2]}$, $\mathcal{L}_{[1,1]}$ and $\mathcal{L}^{\prime}_{[1,1]}$ over $M_{[2]}$, $M_{[1,1]}^{\mathrm{ab}}$ and $M_{[1,1]}^{\mathrm{nab}}$, respectively,
  • then calculate the intersection cohomology groups $IH^{\bullet}(\overline{M}_{[2]},\mathcal{L}_{[2]})$,
    $IH^{\bullet}(\overline{M}_{[1,1]}^{\mathrm{ab}},\mathcal{L}_{[1,1]})$ and $IH^{\bullet}(M_{[1,1]}^{\mathrm{nab}},\mathcal{L}^{\prime}_{[1,1]})$.
The fiber of $\mathcal{L}_{[2]}$ at a point $V\in M_{[2]}$ is simply $H^{\bullet}(\pi^{-1}(V))$, and thus $\mathcal{L}_{[2]}=\mathbb{Q}\oplus\mathbb{Q}[{-2}]$ (cf. §2.3). We thus have
$\mathrm{Hilb}_{-t}(IH^{\bullet}(\overline{M}_{[2]},\mathcal{L}_{[2]}))=(1+t^{2})\cdot IP_{-t}(\mathcal{M}_{0}(2)),$
which is the first term appearing on the right hand side of (9).
Next, we proceed to case (2), the stratum ${M_{[1,1]}^{\mathrm{ab}}}\subset\mathcal{M}_{0}(1)^{2}/\Sigma_{2}$,
${M_{[1,1]}^{\mathrm{ab}}}=\{(L_{1},L_{2})\in\mathcal{M}_{0}(1)^{2}|\,L_{1}\not\cong L_{2}\}/\Sigma_{2}.$
Equality (1.7.B) of the Decomposition Theorem in our situation has the form
$H^{\bullet}(\pi^{-1}(L_{1}\oplus L_{2}))={\mathcal{L}_{[1,1]}}_{L_{1}\oplus L_{2}}\oplus IC(\overline{M}_{[2]},\mathbb{Q})_{L_{1}\oplus L_{2}}.$
Now using formulas (11), (12), we conclude that to compute the stalk of $\mathcal{L}_{[1,1]}$, at $L_{1}\oplus L_{2}\in M_{[1,1]}^{\mathrm{ab}}$, one needs to ”subtract” the $\mathbb{Z}$-graded $\Sigma_{2}$-module
$(\mathbb{Q}\oplus\mathbb{Q}[{-2}])\otimes\mathbb{Q}[x_{1}-x_{2}]/\langle(x_{1}-x_{2})^{g-1}\rangle\text{ from }\mathbb{Q}[x_{1},x_{2}]/\langle x_{1}x_{2},x_{1}^{g},x_{2}^{g}\rangle.$
A simple calculation shows that the difference is 1-dimensional and may be identified with $\mathbb{Q}\cdot(x_{1}-x_{2})^{g-1}\subset\mathbb{Q}[x_{1},x_{2}]/\langle x_{1}x_{2},x_{1}^{g},x_{2}^{g}\rangle.$ This means that
$\mathcal{L}_{[1,1]}=\mathbb{Q}_{\mathcal{M}_{0}(1)^{2}\setminus\Delta}\times_{\Sigma_{2}}\left[\mathbb{S}[{-2}]\right]^{\otimes g-1},$
where $\Delta\subset\mathcal{M}_{0}(1)^{2}$ is the diagonal, and $\mathbb{S}[-2]$ is the 1-dimensional sign representation of $\Sigma_{2}$ put in degree 2. The cohomology ring of $\mathcal{M}_{0}(1)^{2}$ splits into the direct sum
$H^{\bullet}(\mathcal{M}_{0}(1)^{2})=H^{\bullet}(\mathcal{M}_{0}(1)^{2})^{+}\oplus H^{\bullet}(\mathcal{M}_{0}(1)^{2})^{-}$
of its invariant and anti-invariant parts under the action of $\Sigma_{2}$ exchanging the copies of $\mathcal{M}_{0}(1)$. It follows from the basic properties of intermediate extensions that
$IH^{\bullet}(\overline{M}_{[1,1]},\mathcal{L}_{[1,1]})=\begin{cases}H^{\bullet}(\mathcal{M}_{0}(1)^{2})^{+}[2-2g],\text{ if }g\text{ is odd,}\\ H^{\bullet}(\mathcal{M}_{0}(1)^{2})^{-}[2-2g],\text{ if }g\text{ is even,}\end{cases}$
and an elementary calculation shows that the Hilbert series of this cohomology group is given by the second term of (9).
Finally, we show that $\mathcal{L}^{\prime}_{[1,1]}=0$, and thus the contribution of $M_{[1,1]}^{\mathrm{nab}}$ to the right hand side of (9) vanishes. We will give a proof of this statement which is valid for general $r$ in §5, but here, for $r=2$ , we present a simple method based on a dimension count.
According to part 0. of Theorem 1.7, $\mathcal{L}^{\prime}_{[1,1]}$ contributes to $H^{i}(\pi^{-1}(V))$ only if
$i\in[\dim\mathcal{P}_{0}(r)-\dim M_{[1,1]}^{\mathrm{nab}}-r(\pi),\dim\mathcal{P}_{0}(r)-\dim M_{[1,1]}^{\mathrm{nab}}+r(\pi)].$
We now compute this range: we have
$\dim\mathcal{P}_{0}(r)=4g-2,\quad\dim M_{[1,1]}^{\mathrm{nab}}=g,\quad\text{and }r(\pi)=1,$
hence the stalk of $\mathcal{L}^{\prime}_{[1,1]}$ at $V$ should lie in $H^{i}(\pi^{-1}(V))$ for $i\in[3g-3,3g-1]$. Since $22\dim\pi^{-1}(V)=2\dim\mathbb{P}^{g-1}=2g-$ 2, this is not possible for $g\geq 2$, hence $\mathcal{L}^{\prime}_{[1,1]}=0$.
3. Moduli spaces and the parabolic projection map
3.1. Stratification of $\mathcal{M}_{0}(r)$
We begin with introducing the necessary notation. A sequence of arbitrary integers will be denoted, as usual, by $(r_{1},\ldots,r_{k})$; we introduce the notation for a partition of the integer $r$:
$\rho=[r_{1},\ldots,r_{k}]\vdash r,$
if $r_{1}+\ldots+r_{k}=r$, and $r_{1}\geq r_{2}\geq\ldots\geq r_{k}>0$. The length of the partition $\rho$, $k$ in this case, will be denoted by $\mathrm{length}(\rho)$.
In what follows, we will use Theorem 1.2 to identify $\mathcal{M}_{0}(r)$ with the set of isomorphism classes of polystable bundles, and we will denote by $\mathcal{M}_{0}(r)^{\mathrm{st}}$ the subset of stable bundles.
To each polystable $V\in\mathcal{M}_{0}(r)$, we can associate a partition $\rho(V)$ corresponding to the ranks of its direct summands. This induces a decomposition of $\mathcal{M}_{0}(r)$ into a disjoint union of locally closed subsets:
(13) $\mathcal{M}_{0}(r)=\bigsqcup\limits_{\rho\vdash r}M_{\rho},$
where for $\rho=[r_{1},\ldots,r_{k}]\vdash r$,
$M_{\rho}=\{V\cong V_{1}\oplus\ldots\oplus V_{k}|\,V_{j}\in\mathcal{M}_{0}(j)^{\mathrm{st}},\,j=1,\ldots,k\}/\cong.$
We have thus $M_{[r]}=\mathcal{M}_{0}(r)^{\mathrm{st}}$. Given two partitions $\mu,\rho\vdash r$, we will write $\mu>\rho$ if $\rho$ is a refinement of $\mu$; in this case, the closure of $M_{\mu}$ contains $M_{\rho}$.
The decomposition (13) is not a stratification, however, since the automorphism group of a polystable $V\cong V_{1}\oplus\ldots\oplus V_{k}$ depends on whether there are isomorphic stable bundles among its components. This motivates the following refinement of the notion of partition.
A multi-partition of $r$ is a sequence of positive integer pairs
(14) $\overline{\rho}=[(\bar{r}_{1},m_{1}),(\bar{r}_{2},m_{2}),\dots,(\bar{r}_{s},m_{s})],$
satisfying
  • $\bar{r}_{1}\geq\bar{r}_{2}\geq\dots\geq\bar{r}_{s}$,
  • $m_{1}\bar{r}_{1}+\dots+m_{s}\bar{r}_{s}=r$, and
  • $\bar{r}_{i}=\bar{r}_{i+1}\Rightarrow m_{i}\geq m_{i+1}$.
Clearly, each multi-partition $\overline{\rho}$ induces a partition by simply repeating $m_{i}$ times the part $\bar{r}_{i}$, and a partition may be viewed as a multi-partition whose multiplicities are all equal to 1. We introduce the following notation:
[…]
Now, for a multi-partition $\overline{\rho}$ given as in (14), we define
$M_{\overline{\rho}}=\left\{V_{1}^{\oplus m_{1}}\oplus\ldots\oplus V_{s}^{\oplus m_{s}}\in\mathcal{M}_{0}(r)\mid V_{i}\in\mathcal{M}_{0}(\overline{r}_{i})^{\mathrm{st}},V_{i}\not\cong V_{j}\text{ for }1\leq i\neq j\leq s\right\}.$
These sets do define a stratification of $\mathcal{M}_{0}(r)$, and we have
(15) $\mathcal{M}_{0}(r)=\bigsqcup\limits_{\overline{\rho}\Vdash r}M_{\overline{\rho}},\quad\text{and}\quad M_{\rho}=\bigsqcup\limits_{\overline{\rho}\rightsquigarrow\rho}M_{\overline{\rho}}.$
3.2. Local structure of $\mathcal{M}_{0}(r)$
Now we describe the local structure of the singularities of the moduli space $\mathcal{M}_{0}(r)$.
Proposition 3.1.
Let $\overline{\rho}$ be a multi-partition as in (14), and $V\in M_{\overline{\rho}}$ be a polystable vector bundle $V\cong V_{1}^{m_{1}}\oplus\ldots\oplus V_{k}^{m_{k}}$. Then
(16) $\mathrm{Aut}(V)\cong\prod_{i=1}^{k}{\mathrm{G}\mathrm{L}}(m_{i},\mathbb{C}).$
In particular, $\mathrm{Aut}(V)$ is abelian if and only if all multiplicities of $\overline{\rho}$ are equal to 1.
Proof.
See [27, Propositions 4.3 and 4.4]. ∎
Definition 3.2.
We will call a polystable vector bundle $V$ abelian if no two of its direct summands are isomorphic, i.e. if the corresponding multi-partition has multiplicities all equal to 1. This concept induces a decomposition of $M_{\rho}$, $\rho\vdash r$, as well as of $\mathcal{M}_{0}(r)$ into abelian and non-abelian parts:
(17) $\mathcal{M}_{0}(r)=\mathcal{M}_{0}(r)^{{\mathrm{ab}}}\bigsqcup\mathcal{M}_{0}(r)^{{\mathrm{nab}}}\quad\text{and }M_{\rho}=M_{\rho}^{\mathrm{ab}}\bigsqcup M_{\rho}^{\mathrm{nab}}.$
Remark 3.3.
Clearly, we have $\overline{M_{\rho}^{\mathrm{ab}}}=\overline{M}_{\rho}$.
Now we recall some well known results on the infinitesimal structure of the moduli space $\mathcal{M}_{0}(r)$. For the proofs, we refer the reader to [22, 28].
Proposition 3.4.
Let $V\in\mathcal{M}_{0}(r)$. Then the affine GIT quotient
$U_{V}\overset{\mathrm{def}}{=}\mathrm{Ext}^{1}(V,V)\sslash\mathrm{Aut}(V)$
is a local étale neighborhood of $V$ in $\mathcal{M}_{0}(r)$.
Remark 3.5.
Observe that for $t\in\mathbb{C}^{*}$, the element $(t,t,\dots,t)\in\mathrm{Aut}(V)$, thought of as $t\cdot\mathrm{id}\in\prod_{i=1}^{k}{\mathrm{G}\mathrm{L}}(m_{i},\mathbb{C})$, acts trivially on $\mathrm{Ext}^{1}(V,V)$, thus only the action of the quotient $\mathrm{Aut}(V)$ with respect to this one-dimensional subgroup is effective.
Remark 3.6.
Rescaling $\mathrm{Ext}^{1}(V,V)$ by $\lambda\in\mathbb{C}^{*}$ commutes with the action of $\mathrm{Aut}(V)$, and thus descends as an action of $\mathbb{C}^{*}$ on $U_{V}$.
Now let $V\cong V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ be an abelian point.; then
$\mathrm{Ext}^{1}(V,V)\cong\bigoplus_{1\leq i,j\leq k}\mathrm{Ext}^{1}(V_{i},V_{j}),$
and in this case,
  • $\mathrm{Aut}(V)$ is a torus $(\mathbb{C}^{*})^{k}$ with an element $(t_{1},\ldots,t_{k})\in\left(\mathbb{C}^{*}\right)^{k}$ acting diagonally on $V_{1}\oplus\ldots\oplus V_{k}$;
  • this induces the rescaling action by $t_{i}^{-1}t_{j}$ on $\mathrm{Ext}^{1}(V_{i},V_{j})$.
Hence we can describe the local neighborhood $U_{V}$ of $V$ in $\mathcal{M}_{0}(r)$ as follows (see [31, §2]).
Proposition 3.7.
Let $\rho\vdash r$ be a partition of $r$ and let $V\cong V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ be an abelian point of the moduli space $\mathcal{M}_{0}(r)$. Then $U_{V}$ is the product of an affine space and an affine toric variety:
$U_{V}\cong\mathbb{C}^{l}\times\mathcal{C}_{V},$
where $l=\sum_{i=1}^{k}(\mathrm{rk}(V_{i})^{2}(g-1)+1)$ and
(18) $\mathcal{C}_{V}=\bigoplus_{1\leq i\neq j\leq k}\mathrm{Ext}^{1}(V_{i},V_{j})\sslash(\mathbb{C}^{*})^{k}.$
The variety $\mathcal{C}_{V}$ is a cone over a projective toric variety, and is the normal cone of $M_{\rho}^{\mathrm{ab}}$ to $\mathcal{M}_{0}(r)$ at $V$.
3.3. Moduli spaces of parabolic bundles
In this subsection, we recall some basic definitions and facts about parabolic bundles on curves. We will only present a simplified theory, necessary for this paper. For a complete description we refer to [24, 4, 32].
  • Fix a point $p\in{C}$. A parabolic bundle on $C$ is a vector bundle $W$ with a subspace $F_{1}\subset W_{p}$ in the fiber over $p$:
    $W_{p}=F_{2}\supsetneq F_{1}\supsetneq F_{0}=0$
    and parabolic weights $c_{1},c_{2}$ assigned to $F_{2}$ and $F_{1}$ respectively, satisfying the conditions
    $10<c_{1}-c_{2}<$ 1 .
    We set $f_{2}=\mathrm{dim}(F_{1})$ and $\mathrm{rank}(W)-f_{2}=f_{1}=\mathrm{dim}(F_{2}/F_{1})$.
  • The parabolic degree [^3] For technical reasons, we have chosen a sign convention opposite to that in the majority of treatments in the literature. and the parabolic slope of $W$ are defined as
    $\mathrm{pardeg}(W)=\mathrm{deg}(W)-f_{1}c_{1}-f_{2}c_{2};\,\,\,\,\,\mathrm{parslope}(W)=\frac{\mathrm{pardeg}(W)}{\mathrm{rank}(W)}.$
  • A morphism $\phi:W\to W^{\prime}$ of parabolic bundles is a morphism of vector bundles satisfying $\phi_{p}(F_{i})\subset{F^{\prime}_{j-1}}$ if $c_{3-i}<c^{\prime}_{3-j}$. In particular, an endomorphism of a parabolic bundle $W$ is a vector bundle endomorphism preserving the flag $F_{0}\subset F_{1}\subset F_{2}$.
  • Denote by $\mathrm{ParHom}(W,W^{\prime})$ the sheaf of parabolic morphisms from $W$ to $W^{\prime}$. Then there is a short exact sequence of sheaves
    (19) $0\to\mathrm{ParHom}(W,W^{\prime})\to\mathrm{Hom}(W,W^{\prime})\to T_{p}\to{0},$
    where $T_{p}$ is a torsion sheaf supported at $p$. The rank of $T_{p}$ is the number of pairs $(i,j)$, s.t. $c_{i}<c^{\prime}_{j}$, multiplied by $f_{i}f^{\prime}_{j}$ (cf. [4]).
  • Taking cohomology, we obtain the following long exact sequence
    (20) $0\to H^{0}(\mathrm{ParHom}(W,W^{\prime}))\to H^{0}(\mathrm{Hom}(W,W^{\prime}))\to\\ \mathbb{C}^{\mathrm{rk}(T_{p})}\to\mathrm{ParExt}^{1}(W,W^{\prime})\to\mathrm{Ext}^{1}(W,W^{\prime})\to 0.$
If $W^{\prime}\subset{W}$ is a subbundle of $W$, then both $W^{\prime}$ and the quotient $W/W^{\prime}$ inherit a parabolic structure from $W$ in a natural way (cf. [24, Definition 1.7]).
  • A parabolic bundle $W$ is stable of weights $(c_{1},c_{2})$, if any proper subbundle $W^{\prime}\subset{W}$ satisfies $\mathrm{parslope}(W^{\prime})<\mathrm{parslope}(W)$.
Note that the parabolic stability condition depends on the parabolic weights only up to adding the same constant to all weights $c_{i}$, counted with multiplicities. In what follows, we will write
$c=(c_{1},\ldots,c_{1},c_{2},\ldots,c_{2})$
for the sequence with $r$ elements, where each weight $c_{i}$ is repeated $f_{i}$ times. Then without loss of generality, we can assume that the parabolic weights of a rank-$r$ and degree-$d$ bundle belong to the set
[…]
Definition 3.8.
We call a vector $c=(c_{1},\ldots,c_{r})\in\mathbb{R}^{r}$ such that $\sum_{i}c_{i}\in\mathbb{Z}$ regular if for any nontrivial subset $\Psi\subset\{1,2,\ldots,r\}$, we have $\sum_{i\in\Psi}c_{i}\notin\mathbb{Z}$.
For fixed rank $r$, degree $d$ and regular $c\in\Delta_{d}$, Mehta and Seshadri [24] constructed a smooth projective moduli space $\mathcal{P}_{d}(c)$ of stable parabolic bundles whose points are in one-to-one correspondence with the set of isomorphism classes of stable parabolic bundles of weight $c$.
In this paper, we will be interested in two particular moduli spaces of parabolic bundles: set $c_{1}=1/2r$ and $c_{2}=(1-r)/2r$; then $(c_{1},c_{1},\ldots,c_{1},c_{2})\in\Delta_{0}$. Introduce the notation
$\mathcal{P}_{0}(r)\overset{\mathrm{def}}{=}\mathcal{P}_{0}((c_{1},c_{1},\ldots,c_{1},c_{2}))$
for the corresponding moduli space of rank-r degree-0 stable parabolic bundles $W$ with parabolic structure given by a line in the fiber $W_{p}$, and let
$\mathcal{P}_{1}(r)\overset{\mathrm{def}}{=}\mathcal{P}_{1}((1+c_{2},c_{1},\ldots,c_{1}))$
be the moduli space of rank-r degree-1 stable parabolic bundles $W$ with parabolic structure given by the choice of hyperplane in the fiber $W_{p}$.
For a regular $c\in\Delta_{d}$, there exist universal bundles $U$ over $\mathcal{P}_{d}(c)\times C$ endowed with a flag $\mathcal{F}_{1}\subset\mathcal{F}_{2}=U_{p}$, and satisfying the standard tautological properties. In general, such universal bundles $U$ are unique only up to tensoring by the pull-back of a line bundle from $\mathcal{P}_{d}(c)$.
Definition 3.9.
We will say that a universal bundle $U$ over $\mathcal{P}_{0}(r)\times C$ is normalized if the line subbundle $\mathcal{F}_{1}\subset{U}_{p}$ is trivial.
The following statement is straightforward.
Proposition 3.10.
(i) Let $(W,F_{1})$ be a stable parabolic bundle which represents a point in $\mathcal{P}_{0}(r)$. Then the vector bundle $W$ is semistable, and the corresponding map
(21) $\pi:\mathcal{P}_{0}(r)\to\mathcal{M}_{0}(r)$
is a morphism of algebraic varieties.
(ii) Let $(W,F_{1})$ be a stable parabolic bundle which represents a point in $\mathcal{P}_{1}(r)$. Then the vector bundle $W$ is stable, and thus represents a point in $\mathcal{M}_{1}(r)$. Moreover, $\mathcal{P}_{1}(r)$ is isomorphic to the projectivization of a universal bundle $U_{1}(r)_{p}$ over $\mathcal{M}_{1}(r)\times\{p\}$:
$\mathcal{P}_{1}(r)\cong\mathbb{P}(U_{1}(r)_{p})\xrightarrow{\pi_{1}}\mathcal{M}_{1}(r).$
3.4. The tautological Hecke correspondence
One can define (cf. [32, §7]) the tautological Hecke operators between the moduli spaces of parabolic bundles with different degrees and parabolic weights as follows. Given a vector bundle $W$ on $C$ with a flag $0=F_{0}\subsetneq F_{1}\subsetneq F_{2}=W_{p}$ in the fiber over $p\in C$, we consider the associated sheaf of sections $\mathcal{W}$ and define the subsheaf
$\mathcal{W}^{\prime}=\{\gamma\in H^{0}(C,\mathcal{W})\,|\,\gamma(p)\subset F_{1}\}\subset\mathcal{W}.$
Then $\mathcal{W}^{\prime}$ is locally free, and thus defines a vector bundle, which we denote by $W^{\prime}$. We consider the associated morphism of vector bundles $W^{\prime}\to W$ and define the flag $0=F^{\prime}_{0}\subsetneq F^{\prime}_{1}\subsetneq F^{\prime}_{2}=W^{\prime}_{p}$ by setting
$F^{\prime}_{1}=\mathrm{ker}(W^{\prime}_{p}\to F_{1}).$
This construction induces a map, the so-called Hecke operator
$\mathcal{H}:(W,F_{1})\mapsto(W^{\prime},F^{\prime}_{1}).$
Proposition 3.11.
[32, Proposition 7.1] The map $\mathcal{H}$ induces an isomorphism of the moduli spaces
$\mathcal{P}_{d}(\underbrace{c_{1},\ldots,c_{1}}_{f_{1}},\underbrace{c_{2},\ldots,c_{2}}_{f_{2}})\overset{\mathcal{H}}{\cong}\mathcal{P}_{d-f_{1}}(\underbrace{c_{2},\ldots,c_{2}}_{f_{2}},\underbrace{c_{1}-1,\ldots,c_{1}-1}_{f_{1}}).$
Corollary 3.12.
The moduli spaces $\mathcal{P}_{1}(r)$ and $\mathcal{P}_{0}(r)$ defined above are isomorphic.
3.5. Interlude on mixed Hodge structures and the Decomposition Theorem
For most of our arguments, the ”elementary” version of the Decomposition Theorem, Theorem 1.7 suffices, but there is an important technical point, which requires a refinement of this statement: we will need to know that the fibers of the map $\pi$ defined in (21) have pure Hodge structure. Below, we give the necessary definitions and background, but we will be brief, as will we only need a rather trivial part of the theory. We refer the reader to [9] and [8, 5] for more details.
3.5.1. Mixed Hodge structures
The $k$th cohomology of a smooth projective variety $Y$ carries a (pure) Hodge structure of weight $k$:
$H^{k}(Y,\mathbb{C})\cong\bigoplus_{p+q=k}H^{p,q}\text{ with }H^{p,q}=\overline{H^{q,p}}.$
This is equivalent to the datum of a filtration $F^{\bullet}$ on $H^{\bullet}(Y,\mathbb{C})$:
$F^{p}H^{k}(X,\mathbb{C})=\bigoplus_{i\geq p}H^{\dim Y-i,i}\text{ and }H^{p,q}=F^{p}H^{k}(Y,\mathbb{C})\cap\overline{F^{q}H^{k}(Y,\mathbb{C})}.$
The cohomology of a not necessarily projective and not necessarily smooth varieties carry mixed Hodge structure (MHS): given such a variety $Y$, this is a pair $(F^{\bullet},W^{\bullet})$ where
  • $F^{\bullet}$ is a a decreasing filtration on $H^{\bullet}(Y)\otimes\mathbb{C}$, called the Hodge filtration;
  • $W^{\bullet}$ is an increasing filtration on $H^{\bullet}(Y)$, called the Weight filtration, such that $Gr^{W}_{k}:=W_{k}/W_{k-1}\otimes\mathbb{C}$ carries a pure Hodge structure of weight $k$ induced by $F^{\bullet}$.
When $Y$ is projective, we have $W_{k}\supset H^{k}(Y,\mathbb{C})$, and we say that the cohomology of $Y$ is pure if $W_{k-1}(H^{k}(Y,\mathbb{C}))=0$. A key, defining criterion of purity is the following:
Proposition 3.13.
A projective algebraic variety $Y$ has pure cohomology if and only if for a (and thus for any) birational map $\widetilde{Y}\to Y$ from a smooth projective variety $\widetilde{Y}$ to $Y$, the pullback map $H^{\bullet}(Y)\to H^{\bullet}(\widetilde{Y})$ is injective.
Some extensions of this notion:
  • Compactly supported cohomology $H^{\bullet}_{c}(Y)$ has MHS compatible with the MHS on cohomology.
  • MHS may be defined on intersection cohomologies and compactly supported intersection cohomologies as well.
  • Finally, one can extend the notion of MHS to intersection cohomologies with twisted coefficients (see §1.2), replacing the constant sheaf $\mathbb{Q}_{Y}$ with any local system $\mathcal{L}$ supported on an open dense subset of $Y$.
3.5.2. MHS and the Decomposition Theorem
It is a deep statement of Saito’s theory [8, 5, 29] that it is always possible to choose an isomorphism in the Decomposition Theorem 1.7 which respects the Mixed Hodge structures in (1.7.C) and (1.7.B). In particular, in (1.7.B), the MHS on the cohomology of the fiber $f^{-1}(p)$ is the direct sum of the MHS on the intersection cohomologies with twisted coefficients on the right hand side:
(22) $\mathcal{H}^{\bullet}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha})\right)_{p}\cong IH^{\bullet}(U_{\alpha},\mathcal{L}_{\alpha}),\quad U_{\alpha}\text{ is a neighborhood of }p\text{ in }\overline{Y}_{\alpha}.$
We will only need a small, rather trivial part of this theory, which will permit us to conclude that $H^{\bullet}(f^{-1}(p))$ has pure Hodge structure.
The exact statement is as follows:
Proposition 3.14.
Let $f:X\rightarrow Y$ be a proper algebraic map of algebraic varieties as in Theorem 1.7, and let $p\in Y$. Assume that for each stratum $Y_{\alpha}$ containing $p$ in its closure, the intersection cohomology $IH^{\bullet}(U_{\alpha})$ (see (22)) has pure Hodge structure, and that $\mathcal{L}_{\alpha}$ has trivial monodromy. Then the MHS on the cohomology of $H^{\bullet}(f^{-1}(p))$ is pure. More generally, the same conclusion holds if the monodromy of $\mathcal{L}_{\alpha}$ is finite, and the variation of Hodge structure defined by $\mathcal{L}_{\alpha}$ is trivial.
Next, we describe a simple situation, when the intersection cohomology of $IH^{\bullet}(U_{\alpha})$ is pure (see [23, Proposition 2.4]).
Definition 3.15.
A semiprojective variety is a quasi-projective variety $X$ with a $\mathbb{C}^{*}$-action, such that:
  • the fixed point set is proper;
  • For every $p\in X$, the limit $\lim_{t\to 0}t\cdot p$ exists.
Proposition 3.16.
Assume that in the setup above, $U_{\alpha}$ is a semiprojective variety. Then the intersection cohomology $IH^{\bullet}(U_{\alpha})$ of $U_{\alpha}$ has pure Hodge structure.
Now, we observe that the $\mathbb{C}^{*}$ scaling action in Remark 3.6 contracts the étale neighbourhood $U_{V}$ to the origin, and this shows that $U_{V}$ is a semiprojective variety.
Corollary 3.17.
The intersection cohomology $IH^{\bullet}(U_{V})$ of $U_{V}$ introduced in Proposition 3.4 has pure Hodge structure.
3.5.3. Weight polynomials
An important invariant of the mixed Hodge structure on the cohomology of a variety $Y$ is the weight polynomial.
Definition 3.18.
Let $Gr^{W}_{k}(H_{c}^{j}(Y)):=W_{k}(H_{c}^{j}(Y))/W_{k-1}(H_{c}^{j}(Y))\otimes\mathbb{C}$ be the weight-$k$ part of the $j$th compactly supported cohomology of $Y$. The weight polynomial of $Y$ is the generating function
$E_{t}(Y)=\sum_{j,k}(-1)^{j}\dim Gr^{W}_{k}(H_{c}^{j}(Y))t^{k}.$
Clearly, if the cohomology of $Y$ is pure, then $E_{t}(Y)$ is the (compactly supported) Poincaré polynomial of $Y$. A key property of the weight polynomial is that it is motivic.
Proposition 3.19.
If $Z\subset Y$ is a locally closed subvariety of the variety $Y$, then
$E_{t}(Y)=E_{t}(Z)+E_{t}(Y\setminus Z).$
3.6. Our plan
Having laid the groundwork, we can now spell out our plan in more detail. We will apply our strategy (see §1.7) to the map $\pi$ introduced in (21). In (15), we described a stratification of $\mathcal{M}_{0}(r)$, which is adapted to the map $\pi$. The Decomposition Theorem (1.7.A), in our case, will take the form
(23) $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}\cong\bigoplus_{\overline{\rho}\Vdash r}IC(\overline{M}_{\overline{\rho}},\mathcal{L}_{\overline{\rho}}),$
where $\mathcal{L}_{\overline{\rho}}$ is a local system on the stratum $M_{\overline{\rho}}$.
We will attack step i. of the strategy in §4 below. We will describe the cohomology of the fibers of the map $\pi$ in the abelian case both by generators and relations (Theorem 4.14), and as graded vector spaces (Theorem 4.12). The Betti numbers of the IC sheaf stalks were calculated in [31].
Then we turn to step ii. of the strategy: the calculation of the local systems $\mathcal{L}_{\overline{\rho}}$, which is the central part of our work. In §5, we show that $\mathcal{L}_{\overline{\rho}}=0$ whenever the multipartition $\overline{\rho}$ is not abelian.
At this point, in order to write down (1.7.B) for our case, and to streamline our notation for the rest of the paper, we introduce the following conventions:
  • we will denote by $\mathcal{L}_{\rho}$ the local system over $M_{\rho}^{\mathrm{ab}}$;
  • for a generic point $V\in M_{\rho}^{\mathrm{ab}}$, we will denote by $F_{\rho}$ the fiber $\pi^{-1}(V)$, and by $L_{\rho}$ the stalk of $\mathcal{L}_{\rho}$ over $V$; we have thus $L_{\rho}\subset H^{\bullet}(F_{\rho})$.
Recall that $\overline{M_{\rho}^{\mathrm{ab}}}=\overline{M}_{\rho}$ (see Remark 3.3); with these preparations, we can write (1.7.A) as
(24) $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}\cong\bigoplus_{\rho\vdash r}IC(\overline{M}_{\rho},\mathcal{L}_{\rho}).$
and (1.7.B) as
(25) $H^{\bullet}(F_{\rho})\cong L_{\rho}\oplus\bigoplus_{\mu>\rho}\mathcal{H}^{\bullet}\left(IC(\overline{M}_{\mu},\mathcal{L}_{\mu})\right)_{V}.$
We present two methods for the calculation of $\mathcal{L}_{\rho}$ in the abelian case:
  • We calculate the Hilbert function of the graded vector space $L_{\rho}$ (Theorem 6.1) via the subtraction method in §6, using ideas from [31]. This is not quite sufficient for our purposes, since we need to know the monodromy of the local system $\mathcal{L}_{\rho}$.
  • In §7, we develop a calculus of the refined intersection forms, which give us detailed information about $L_{\rho}$. In particular, it shows that a certain concrete element $w_{\rho}\in F_{\rho}$ is in $L_{\rho}$ (Theorem 7.3).
Probably, either of these methods can be pushed through to identify $L_{\rho}$ inside $H^{\bullet}(F_{\rho})$, but, as it happens, the two results we obtain, combined with the fact that $L_{\rho}\subset H^{\bullet}(F_{\rho})$ is an ideal, lead to an elegant proof of our description of $\mathcal{L}_{\rho}$, Theorem 7.7.
At this point, we will be ready for the final step iii.; the equality (1.7.C) has the following form in our case:
$H^{\bullet}(\mathcal{P}_{0}(r))\cong\bigoplus_{\rho\vdash r}IH^{\bullet}(\overline{M}_{\rho},\mathcal{L}_{\rho}).$
Proposition 3.10 and Corollary 3.12 tell us that (as graded vector spaces). Finally, a recursion leading to our main result, Theorem 1.4 is developed in §8.
4. Cohomology rings of the fibers of the map $\pi$
We showed in Proposition 3.10 that the forgetful map $\pi:\mathcal{P}_{0}(r)\to\mathcal{M}_{0}(r)$, which sends a stable parabolic bundle $(W,l\subset W_{p})$ to the underlying semistable bundle $W$, is proper. In this section, we analyze the cohomology of the fibers of this map over the abelian points of our moduli space $\mathcal{M}_{0}(r)$. The main results are Theorem 4.12, where we calculate the Betti numbers of the fiber, and Theorem 4.14, where we describe its cohomology ring via generators and relations.
4.1. Towers of projective spaces
Let $(r_{1},r_{2},\dots,r_{k})$ be a sequence of positive integers whose sum is $r$. Given a sequence of $k$ pairwise non-isomorphic, stable, degree-0 vector bundles $(V_{1},V_{2},\ldots,V_{k})$ with $\mathrm{rank}(V_{i})=r_{i},\,i=1,\dots,k$, we define
$\mathcal{T}_{(V_{1},\ldots,V_{k})}\overset{\mathrm{def}}{=}\{(W,l\subset W_{p})\in\mathcal{P}_{0}(r),\;\text{s.t. }\;\exists\;W=W_{1}\supset W_{2}\supset\ldots\supset W_{k+1}=\{0\}\\ \text{ with }W_{j}/W_{j+1}\cong V_{j}\;\text{for}\;j=1,\ldots,k\}\subset\mathcal{P}_{0}(r).$
Proposition 4.1.
The space $\mathcal{T}_{(V_{1},\ldots,V_{k})}$ defined above is a smooth projective variety of dimension $\sum_{i<j}r_{i}r_{j}(g-1)+r-k$, which is an iterated fibration by projective spaces.
Remark 4.2.
We will call such spaces, informally, towers.
Proof.
We start by describing the parabolic structure on the quotients $V_{1},\ldots,V_{k}$. Let $(W,l)\in\mathcal{T}_{(V_{1},\ldots,V_{k})}$. Note that for $j>1$:
  • the quotients $W/W_{j}$ (in particular, $W/W_{2}\cong V_{1}$) inherit a parabolic structure from $W$, which is given by the image of the line $l\subset W_{p}$, and parabolic weights $c_{1}=1/2r,c_{2}=(1-r)/2r$;
  • the line $l$ is not contained in $W_{j|p}$, and thus we will treat the subbundle $W_{j}\subset W$ as a degree-0 parabolic subbundle of $W$ with trivial parabolic structure and parabolic weights all equal to $c_{1}=1/2r$;
  • as a result, the quotient $W_{j}/W_{j+1}\cong V_{j}$ inherits from $W_{j}$ a trivial parabolic structure with parabolic weights all equal to $c_{1}$.
In what follows, we consider these bundles as parabolic bundles with parabolic structures described above.
We prove the proposition by induction on $k$. First, note that the parabolic structures on the quotients $W/W_{2}\cong V_{1}$ are parametrized by the points of the projective space $\mathcal{T}_{(V_{1})}\cong\mathbb{P}(V_{1|p})$, and, over $\mathbb{P}(V_{1|p})\times C$, there is a universal bundle
(26) $\mathcal{W}^{(1)}\cong\mathcal{O}(1)\boxtimes V_{1}\to\mathbb{P}(V_{1|p})\times C,$
such that for every $l\in\mathbb{P}(V_{1|p})$, the restriction of $\mathcal{W}^{(1)}$ to $\{l\}\times C$ represents the vector bundle $V_{1}$ with parabolic structure given by $l$, and $\mathcal{W}^{(1)}$ is normalized in such way that $\mathcal{W}^{(1)}|_{\mathcal{T}_{(V_{1})}\times\{p\}}$ has a non-vanishing tautological section
$\mathcal{O}\hookrightarrow\mathcal{W}^{(1)}|_{\mathcal{T}_{(V_{1})}\times\{p\}}.$
Now for $i\geq 1$ we denote by $\underline{i}$ the sequence $(V_{1},V_{2},\ldots,V_{i})$, and let $\mathcal{W}^{\underline{i}}\to\mathcal{T}_{(V_{1},\ldots,V_{i})}\times C$ be the normalized universal bundle, such that for every point $w\in\mathcal{T}_{(V_{1},\ldots,V_{i})}$ the restriction of $\mathcal{W}^{\underline{i}}$ to $\{w\}\times C$ represents the vector bundle given by $w$. With a slight abuse of notation, for every $i\geq 1$ we will denote by the same symbols $\xi$ and $\zeta$ the projections
[…]
We define the sheaf of morphisms
$\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})=\{\varphi:\mathcal{W}^{\underline{i}}\to\zeta^{*}V_{i+1}\;|\\ \forall w\in{\mathcal{T}}_{(V_{1},\ldots,V_{i})}\,\,\,\varphi|_{\{w\}\times C}\text{\, is a morphism of parabolic bundles}\}.$
over $\mathcal{T}_{(V_{1},\ldots,V_{i})}\times C$. A simple calculation shows that $\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})$ is a subsheaf of $\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})$ with quotient isomorphic to $\zeta^{*}V_{i+1}|_{{{\mathcal{T}}_{(V_{1},\ldots,V_{i})}\times\{p\}}}$:
(27) $0\to\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to\zeta^{*}V_{i+1}|_{{{\mathcal{T}}_{(V_{1},\ldots,V_{i})}\times\{p\}}}\to 0,$
which leads to the long exact sequence
(28) $0\to R^{0}\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to\zeta^{*}V_{i+1}|_{{{P}_{(V_{1},\ldots,V_{i})}\times\{p\}}}\to\\ \to R^{1}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to R^{1}\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to 0.$
Now the inductive step may be cast in the following form.
Lemma 4.3.
Let $i\geq 1$.
(i) The set $\mathcal{T}_{(V_{1},\ldots,V_{i+1})}$ is in bijection with the total space of the projective bundle
$\mathbb{P}(R^{1}\zeta_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\xi^{*}V_{i+1}))\xrightarrow{\psi_{i}}\mathcal{T}_{(V_{1},\ldots,V_{i})}.$
(ii) On $\mathcal{T}_{(V_{1},\ldots,V_{i+1})}\times C$, there exists a universal family of extensions
$0\to\xi^{*}\mathcal{O}(1)\otimes\zeta^{*}V_{i+1}\to\mathcal{W}^{\underline{i+1}}\to(\psi_{i}\times\mathrm{id})^{*}\mathcal{W}^{\underline{i}}\to 0,$
such that for every point $w\in\mathcal{T}_{(V_{1},\ldots,V_{i+1})}$ the restriction of $\mathcal{W}^{\underline{i+1}}$ on $\{w\}\times C$ represents the class of extension given by $w$.
(iii) The normalized universal bundle on $\mathcal{P}_{0}(r_{1}+\ldots+r_{i+1})\times C$ (cf. notation on page 3.9) restricts to the universal bundle $\mathcal{W}^{\underline{i+1}}$ on $\mathcal{T}_{(V_{1},\ldots,V_{i+1})}\times C$.
Proof.
Part (i) is trivial. For (ii) note that $R^{0}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})=0$ and use [21]. Part (iii) follows from the construction of $\mathcal{W}^{\underline{i+1}}$. ∎
We thus conclude that the space $\mathcal{T}_{(V_{1},\ldots,V_{k})}$ is a tower, i.e. an iterated fibration of projective spaces:
(29) $\mathbb{P}(V_{1|p})\xleftarrow{\psi_{1}}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{1}},\zeta^{*}V_{2}))\xleftarrow{\psi_{2}}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{2}},\zeta^{*}V_{3}))\xleftarrow{\psi_{3}}\ldots\\ \ldots\xleftarrow{\psi_{k-2}}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{k-2}},\zeta^{*}V_{k-1}))\xleftarrow{\psi_{k-1}}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{k-1}},\zeta^{*}V_{k})).$
To calculate its dimension, observe that for $j=2,\ldots,k$ the fiber dimension of $\psi_{j-1}$ is equal to
$\mathrm{dim}\mathbb{P}(\mathrm{ParExt}^{1}(W/W_{j},V_{j}))=\mathrm{dim}H^{1}(\mathrm{Hom}(W/W_{j},V_{j}))-\\ \mathrm{dim}H^{0}(\mathrm{Hom}(W/W_{j},V_{j}))+\mathrm{dim}H^{0}(\mathrm{ParHom}(W/W_{j},V_{j}))+r_{j}-1=\\ r_{j}(r_{1}+r_{2}+\ldots+r_{j-1})(g-1)+r_{j}-1.$
To obtain the first equality we applied the long exact sequence (20) to the pair of parabolic bundles $W/W_{j}$ and $V_{j}$; for the last equality we used that $\mathrm{parslope}(W/W_{j})<\mathrm{parslope}(V_{j})$, and thus the sheaf $\mathrm{ParHom}(W/W_{j},V_{j})$ has no nontrivial global sections (cf. [30, Proposition 6]).
Finally, we note that
$r_{1}-1+\sum_{j=2}^{k}(r_{j}(r_{1}+r_{2}+\ldots+r_{j-1})(g-1)+r_{j}-1)=\sum_{i<j}r_{i}r_{j}(g-1)+r-k,$
and the proof of Proposition 4.1 is complete. ∎
Corollary 4.4.
In the course of the proof, we showed that
(30) $P_{t}(\mathcal{T}_{(V_{1},\ldots,V_{k})})=\prod_{i=1}^{k}\frac{t^{2r_{i}(r_{1}+\ldots+r_{i-1})+2r_{i}}-1}{t^{2}-1}.$
4.2. The cohomology ring of the projective tower
The following proposition describes the cohomology ring of the projective tower (29).
Proposition 4.5.
The cohomology ring $H^{\bullet}(\mathcal{T}_{(V_{1},\ldots,V_{k})})$ is isomorphic to the quotient of $\mathbb{Q}[x_{1},\ldots,x_{k}]$ by the ideal $I_{(r_{1},\ldots,r_{k})}$ generated by the polynomials
$\phi_{1}(x_{1},\ldots,x_{k})=x_{1}^{r_{1}}\text{\,\,\, and\,\,\,}\phi_{i}(x_{1},\ldots,x_{k})=x_{i}^{r_{i}}\cdot\prod_{j=1}^{i-1}(x_{i}-x_{j})^{r_{i}r_{j}(g-1)}\text{\,\, for\,\,\,}2\leq i\leq k.$
Proof.
It is easy to see that $\mathbb{Q}[x_{1},\ldots,x_{k}]/I_{(r_{1},\ldots,r_{k})}$ has the Hilbert function (30), thus it is sufficient to identify the relations $\phi_{i}$, $i=1,\dots,k$, in the cohomology of our tower.
Introduce the notation $x_{1}=c_{1}(\mathcal{O}(1)\to\mathbb{P}(V_{1|p}))$, and
(31) $x_{i}=c_{1}(\mathcal{O}(1)\to\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i-1}},\zeta^{*}V_{i})))\text{\,\, for \,\,}i=2,\ldots,k.$
Clearly, we have $x_{1}^{r_{1}}=0$.
Putting equation (26) and Lemma 4.3 (ii) together, we conclude the following.
Lemma 4.6.
The universal bundle $\mathcal{W}^{\underline{i}}$, $1\leq i\leq k$, has Chern roots $\xi^{*}x_{j}+\zeta^{*}\alpha^{j}_{l}$, $j=1,2,\ldots,i$, $l=1,\ldots,r_{j}$, where $\alpha^{j}_{l}\in H^{2}(C)$ are the Chern roots of the vector bundle $V_{j}$.
Now we calculate the Chern roots of the vector bundle
(32) $R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})\to\mathcal{T}_{(V_{1},\ldots,V_{i})}.$
First, note that since $R^{0}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})=0$, it follows from the short exact sequence (27) that
$ch(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1}))=-ch(\xi_{!}\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1}))+r_{i+1}.$
Next, applying the Grothendieck-Riemann-Roch theorem we obtain
(33) $-ch(\xi_{!}\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1}))=-\xi_{*}(ch(\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1}))(1-(g-1)\omega))=\\ (g-1)ch(\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1}))|_{{{P}_{(V_{1},\ldots,V_{i})}\times\{p\}}}+\xi_{*}ch(\mathrm{Hom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})),$
where $\omega\in H^{2}(C)$ is the class of the curve. Using Lemma 4.6, we arrive at the following result.
Lemma 4.7.
The vector bundle (32) has Chern roots:
  • $-x_{j}\text{ with multiplicity }r_{j}r_{i+1}(g-1)\text{ for }j=1,\ldots,i$;
  • 0 with multiplicity $0\text{ with multiplicity }r_{i+1}$.
Now, taking the projectivization of (32), we arrive at the relation
$x_{i+1}^{r_{i+1}}\prod_{j=1}^{i}(x_{i+1}-x_{j})^{r_{j}r_{i+1}(g-1)}=0$
in the cohomology ring $H^{\bullet}(\mathcal{T}_{(V_{1},\ldots,V_{k})})$. ∎
Following the same argument, one finds an integral formula for the tautological classes over our tower.
Corollary 4.8.
Consider the variables $x_{1},\dots,x_{k}$ as elements of the second cohomology of the tower $\mathcal{T}_{(V_{1},\ldots,V_{k})}$, as described above. Then, for any polynomial $\phi\in\mathbb{Q}[x_{1},\ldots,x_{k}]$,
(34) $\operatornamewithlimits{\int}_{\mathcal{T}_{(V_{1},\ldots,V_{k})}}\phi(x_{1},\dots,x_{k})=\\ \operatornamewithlimits{\mathrm{IRes}}_{x_{1},\dots,x_{k}=0}\frac{\phi(x_{1},\dots,x_{k})\;dx_{1}\dots dx_{k}}{x_{1}^{r_{1}}\cdot x_{2}^{r_{2}}(x_{2}-x_{1})^{r_{2}r_{1}(g-1)}\cdot\ldots\cdot x_{k}^{r_{k}}\prod_{i<k}(x_{k}-x_{i})^{r_{k}r_{i}(g-1)}},$
where the iterated residue means first taking the ordinary one-variable residue with respect to $x_{k}$, while keeping the other variables as generic constants, and then iterating this operation. In terms of residue integrals we have:
$\operatornamewithlimits{\mathrm{IRes}}_{x_{1},\dots,x_{k}=0}\psi(x_{1},\dots,x_{k})=\int_{\{|x_{i}|=\varepsilon_{i}\}}\psi(x_{1},\dots,x_{k}),\text{ where }0<\varepsilon_{1}\ll\varepsilon_{2}\ll\dots\ll\varepsilon_{k}.$
A similar calculation shows the following formula.
Proposition 4.9.
The Chern character of the tangent bundle of $\mathcal{T}_{(V_{1},\ldots,V_{k})}$ is equal to
$ch(T\mathcal{T}_{(V_{1},\ldots,V_{k})})=\sum_{1\leq i<j\leq k}r_{i}r_{j}(g-1)\exp({x_{j}-x_{i}})+\sum_{1\leq i\leq k}r_{i}\exp({x_{i}})-k.$
4.3. The Poincaré polynomial of the abelian fiber
Now we begin the study of the cohomology of the fibers of the map $\pi$. Fix a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$. Recall the notation $F_{V}:=\pi^{-1}(V)$ for the fiber of the map $\pi$.
The following statement is straightforward.
Proposition 4.10.
  1. (i)
    Suppose $V\cong V_{1}\oplus\ldots\oplus V_{k}$ represents a point in $M_{\rho}^{\mathrm{ab}}$. Then the fiber $F_{V}$ is a union of projective towers:
    $F_{V}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})},$
  2. (ii)
    and
    $\bigcap_{\begin{subarray}{c}\sigma\in\Sigma_{k}\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}\cong\prod_{i=1}^{k}\mathbb{P}(V_{i|p}).$
Our next goal is to provide a recursive formula for the Betti numbers of the abelian fibers. We will need the following statement.
Lemma 4.11.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, $V\in M_{\rho}^{\mathrm{ab}}$. Then
  1. (i)
    the weight polynomial (see §3.5.3) of the fiber $F_{V}$ coincides with its Poincaré polynomial: $E_{t}(F_{V})=P_{t}(F_{V})$, and
  2. (ii)
    the odd cohomologies of $F_{V}$ vanish.
Proof.
Our starting point is the representation of the cohomology of $F_{\rho}$ via the Decomposition Theorem (see (25)). It is easy to verify that abelian strata are in the closure of abelian strata only. As abelian strata have no local monodromy, Proposition 3.14, and Corollary 3.17 imply that the cohomology of $F_{V}$ is pure. This means, in particular, that that the weight polynomial of $F_{V}$ coincides with its Poincaré polynomial: $E_{t}(F_{V})=P_{t}(F_{V})$.
According to Proposition 4.10, we have a surjective birational map from a disjoint union of smooth towers to $F_{V}$:
$\sqcup_{\sigma\in\Sigma_{k}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}\to F_{V}.$
Now using Proposition 3.13, we can conclude that the map
(35) $H^{\bullet}(F_{V})\hookrightarrow\oplus_{\sigma\in\Sigma_{k}}H(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$
is an injection, and thus (see Proposition 4.5) the fiber $F_{V}$ has only even cohomology. ∎
Theorem 4.12.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, $V\in M_{\rho}^{\mathrm{ab}}$ and let $f(\rho;t)$ be the polynomial such that $f(\rho;t^{2})=P_{t}(F_{V})$. For any subset $J=\{j_{1},j_{2},\ldots,j_{s}\}\subset\{1,2,\ldots,k\}$, denote by $\overline{J}=\{1,\ldots,k\}\setminus J$ its complement, and by $\rho_{J}=[r_{j_{1}},\ldots,r_{j_{s}}]$ the corresponding partition of $r_{J}=\sum_{j\in J}r_{j}$. Setting $f(\rho_{\emptyset};t)=1$, and $r_{\emptyset}=0$, the recursive formula
(36) $f(\rho;t)=\sum_{J\subset\{1,\ldots,k\}}(-1)^{|J|-1}f(\rho_{\overline{J}};t)\prod_{j\in J}p(r_{j}r_{\overline{J}}(g-1)+r_{j};t)$
holds, where $p(n;t)=(t^{n}-1)/(t-1)$, and the sum is taken over all non-empty subsets $J\subset\{1,2,\ldots,k\}$.
Proof.
Let $V=V_{1}\oplus\dots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$, and, for $\,i=1,\ldots,k$, consider the subset of $F_{V}$ defined as
$\mathcal{F}_{i}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\\ \sigma(k)=i\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}.$
Note that $\mathcal{F}_{i}$, is the set of those elements of the fiber $F_{V}$ which have $V_{i}$ as a subbundle. It follows then, that for a subset $J\subset\underline{k}$, the set
$\mathcal{F}_{J}=\bigcap_{i\in J}\mathcal{F}_{i}$
is the set of bundles in the fiber $F_{V}$ which have $\oplus_{i\in J}V_{i}$ as a subbundle.
According to Lemma 4.3 (i), $\mathcal{F}_{i}$ is a projective bundle over a ”smaller” fiber:
(37) $\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{V/V_{i}},\zeta^{*}V_{i}))\to F_{V/V_{i}},$
where following the notation from §4.1, we denoted by $\mathcal{W}^{V/V_{i}}$ the normalized universal bundle over $F_{V/V_{i}}\times C$, and by $\xi$ and $\zeta$ the projections along $C$ and $F_{V/V_{i}}$, respectively.
Since $\mathcal{F}_{i}$ is a projective bundle (37) with fiber dimension $r_{i}(r-r_{i})(g-1)+r_{i}-1$ over $F_{V/V_{i}}$, its cohomology is pure (cf. Lemma 4.11(i)), and thus we can identify the term in (36) corresponding to $J=\{i\}$ as the weight polynomial of $\mathcal{F}_{i}$.
Similarly, the set $\mathcal{F}_{ij}$ can be identified with the bundle
$\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{V/(V_{i}\oplus V_{j})},\zeta^{*}V_{i}))\times\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{V/(V_{i}\oplus V_{j})},\zeta^{*}V_{j}))\to\\ F_{V/(V_{i}\oplus V_{j})},$
and thus the term in (36) corresponding to $J=\{i,j\}$ may be interpreted as the weight polynomial of $\mathcal{F}_{ij}$.
Now observe that $\{\mathcal{F}_{i},\,i=1,\dots,k\}$ is a covering of $F_{V}$, and the corresponding inclusion-exclusion formula results in the identity
$F_{V}=\sum_{i=1}^{k}\mathcal{F}_{i}-\sum_{i<j}\mathcal{F}_{ij}+\sum_{i<j<m}\mathcal{F}_{ijm}-\dots$
in the Grothendieck ring of varieties, which, using Proposition 3.19, implies the equality of the weight polynomials
(38) $E_{t}(F_{V})=\sum_{i=1}^{k}E_{t}(\mathcal{F}_{i})-\sum_{i<j}E_{t}(\mathcal{F}_{ij})+\sum_{i<j<m}E_{t}(\mathcal{F}_{ijm})-\dots.$
This is exactly the equality (36) (cf. Lemma 4.11), and this completes the proof of Theorem 4.12. ∎
4.4. Intersection of projective towers
Our next project is to study the intersection between two towers, which will be key in §7. Our construction is a generalization of the tower construction of §4.1. To describe this generalization, we will need to introduce a combinatorial device associated to a pair of permutations: $\sigma,\tau\in\Sigma_{k}$.
Introduce the partial order $\prec$ on $\underline{k}=\{1,\ldots,k\}$:
$j\prec i\quad\text{if}\quad\tau^{-1}(j)<\tau^{-1}(i)\text{ and }\sigma^{-1}(j)<\sigma^{-1}(i).$
Using this partial order, we construct and ordered decomposition of the set $\underline{k}$
$\Pi^{\sigma\tau}=(\Pi_{1},\dots,\Pi_{n}),\quad\underline{k}=\Pi_{1}\sqcup\dots\sqcup\Pi_{n}$
as follows.
The first element of the partition $\Pi_{1}\subset\underline{k}$ is the set of minimal elements of $\underline{k}$ with respect to $\prec$. Next, $\Pi_{2}$ is the set of minimal elements in $\underline{k}\setminus\Pi_{1}$. One continues this process, defining the set $\Pi_{m}$ by
$\Pi_{m}=\{i\in\underline{k}\setminus(\Pi_{1}\cup\dots\cup\Pi_{m-1})|\;j\prec i\Rightarrow j\in\Pi_{1}\cup\dots\cup\Pi_{m-1}\}.$
Clearly, this process eventually exhausts all of $\underline{k}$.
Now let $V\cong V_{1}\oplus\ldots\oplus V_{k}$ be a point in $M_{\rho}^{\mathrm{ab}}$. With these preparations, we can mimic the construction of Lemma 4.3.
  • We start with $\Pi_{1}$, and to each number $i\in\Pi_{1}$, we associate the projective space $\mathbb{P}(V_{i|p})$ together with the normalized universal bundle (cf. (26))
    $\mathcal{W}_{i}=\mathcal{O}(1)\boxtimes V_{i}\to\mathbb{P}(V_{i|p})\times C,$
    and we set
    $\mathcal{I}(\sigma,\tau)^{\leq{1}}=\prod_{i\in\Pi_{1}}\mathbb{P}(V_{i|p}),$
    Now we iterate the following procedure, until we arrive at $\Pi_{n}$:
  • Let $i\in\Pi_{m}$ and let $\{j_{1},\ldots,j_{s}\}=\{j\in\Pi_{m-1}|\,j\prec i\}$. Then to $i$, we associate the projective bundle
    (39) $\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}_{j_{1}}\oplus\ldots\oplus\mathcal{W}_{j_{s}},\zeta^{*}V_{i}))\to\mathcal{I}(\sigma,\tau)^{\leq{m-1}}$
    together with the normalized universal bundle (cf. notation on page 3.9 and Lemma 4.3(iii))
    $\mathcal{W}_{i}\to\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}_{j_{1}}\oplus\ldots\oplus\mathcal{W}_{j_{s}},\zeta^{*}V_{i}))\times C.$
  • We set $\mathcal{I}(\sigma,\tau)^{\leq{m}}$ to be the total space of the bundle
    $\prod_{i\in\Pi_{m}}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}_{j_{1}}\oplus\ldots\oplus\mathcal{W}_{j_{s}},\zeta^{*}V_{i}))\to\mathcal{I}(\sigma,\tau)^{\leq{m-1}}.$
  • Finally, we set $\mathcal{I}(\sigma,\tau)\overset{\mathrm{def}}{=}\mathcal{I}(\sigma,\tau)^{\leq{n}}$.
An example is displayed on Figure 2. The edges are drawn between comparable pairs: $j\prec i$ from left to right.
2222111133334444
${\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}_{2}\oplus\mathcal{W}_{1},\zeta^{*}V_{3}))\times\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}_{2},\zeta^{*}V_{4}))}$${\mathbb{P}(V_{1|p})\times\mathbb{P}(V_{2|p})}$
Figure 2. The construction for $\sigma=(1243)$ and $\tau=\mathrm{id}$; here $\Pi_{1}=\{1,2\}$ and $\Pi_{2}=\{3,4\}$.
Proposition 4.13.
Let $V\cong V_{1}\oplus\ldots\oplus V_{k}$ be a point in $M_{\rho}^{\mathrm{ab}}$. The intersection of the two towers of projective bundles
$\mathcal{T}_{\tau}=\mathcal{T}_{(V_{\tau(1)},\ldots,V_{\tau(k)})}\text{\,\, and\,\, }\mathcal{T}_{\sigma}=\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}$
is isomorphic to the space constructed above:
$\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}\cong{{\mathcal{I}(\sigma,\tau).}}$
Proof.
Comparing our construction of $\mathcal{I}(\sigma,\tau)$ to Lemma 4.3(i), we note that at every step, the projective bundle (39) associated to the vertex $i$ is a subbundle of the projective bundle on the ${\sigma^{-1}(i)}^{\mathrm{th}}$ level in the tower $\mathcal{T}_{\sigma}$. The same is true for $\mathcal{T}_{\tau}$, and thus $\mathcal{I}(\sigma,\tau)\subset\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}$.
Now to any parabolic bundle $W\in\mathcal{T}_{\sigma}$:
(40) $W=W_{1}\supset W_{2}\supset\ldots\supset W_{k}\supset W_{k+1}=0\text{\, with \, }W_{j}/W_{j+1}\cong V_{\sigma(j)},\text{\, for \, }j=1,\ldots,k.$
we can associate a partial order $\prec_{W}$ on $\underline{k}$ compatible with the total order induced by $\sigma$:
$\sigma(j)\prec_{W}\sigma(i)\quad\text{if}\quad W_{j}\supsetneq W_{i}\;\text{and}\;W_{j}/W_{i+1}\not\supset V_{\sigma(i)}\oplus V_{\sigma(j)}.$
Clearly, this happens only if $j<i$. In other words, for any $W\in\mathcal{T}_{\sigma}$ we obtain that $j\prec_{W}i$ implies $\sigma^{-1}(j)<\sigma^{-1}(i)$. Hence if $W\in\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}$ then the partial order $\prec_{W}$ has to be compatible with the partial order
$j\prec i\text{\; if \;}\sigma^{-1}(j)<\sigma^{-1}(i)\text{\; and \;}\tau^{-1}(j)<\tau^{-1}(i),$
and thus $W$ have to be an element of $\mathcal{I}(\sigma,\tau)$; this completes our proof. ∎
4.5. The cohomology ring of the abelian fiber
Now we are ready to describe the cohomology ring of the fiber for any point $V$ in $\mathcal{M}_{0}(r)^{\mathrm{ab}}$.
Theorem 4.14.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, and let $V$ be a point in $M_{\rho}^{\mathrm{ab}}$. The cohomology ring $H^{\bullet}(F_{V})$ is generated by $k$ elements $x_{1},\ldots,x_{k}\in H^{2}(F_{V})$ with the only relations
(41) $\phi_{D}(x_{1},\ldots,x_{k})=\prod_{i\in A}x_{i}^{r_{i}}\cdot\prod_{\begin{subarray}{c}i\in A,j\in B\end{subarray}}(x_{i}-x_{j})^{r_{i}r_{j}(g-1)},$
where $D=A\sqcup B$ runs over all partitions of the set $\{1,2,\ldots,k\}$, such that $A\neq\emptyset$.
We start with defining the elements $x_{1},\ldots,x_{k}\in H^{2}(F_{V})$ appearing in the theorem. For any $\sigma\in\Sigma_{k}$, let
$x^{\sigma}_{\sigma(1)}=c_{1}(\mathcal{O}(1)\to\mathbb{P}(V_{\sigma(1)|p}))\text{\,\, and \,\,}\\ x^{\sigma}_{\sigma(i)}=c_{1}(\mathcal{O}(1)\to\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{(V_{\sigma(1)},\ldots,V_{\sigma(i-1)})},\zeta^{*}V_{\sigma(i)}))),\,\,\,\,2\leq i\leq k$
be the degree-2 generators of $H^{\bullet}(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$ (cf. notation in the proof of Proposition 4.5).
Lemma 4.15.
There exist $k$ elements $x_{1},x_{2},\ldots,x_{k}\in H^{\bullet}(F_{V})$, such that for any $1\leq i\leq k$ the restriction of $x_{i}$ to ${\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}}$ coincides with $x^{\sigma}_{i}$.
Proof.
To prove the statement, we need to show that for any $1\leq i\leq k$ and any permutations $\sigma,\tau\in\Sigma_{k}$ the restrictions of the cohomology classes $x^{\sigma}_{i}$ and $x^{\tau}_{i}$ to
$\mathcal{I}(\sigma,\tau)={\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}}\cap{\mathcal{T}_{(V_{\tau(1)},\ldots,V_{\tau(k)})}}$
coincide. Repeating the argument from the proof of Proposition 4.13 we note that the projective bundle (39) is a subbundle of the projective bundle on the $\sigma^{-1}(i)^{\mathrm{th}}$ level in the tower $\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}$, and thus $x^{\sigma}_{i}$ restricts to the first Chern class of the projective bundle (39). Since the same holds for the cohomology calss $x^{\tau}_{i}$, the result follows. ∎
Now we show that the relations (41) are satisfied in $H^{\bullet}(F_{V})$.
Proposition 4.16.
Let $x_{1},\ldots,x_{k}$ be the classes in $H^{2}(F_{V})$ defined in Lemma 4.15. Then for any partition $D=A\sqcup B$ of $\{1,2,\ldots,k\}$, such that $A\neq\emptyset$ the polynomial $\phi_{D}(x_{1},\ldots,x_{k})$ defined in (41) is a zero element in the cohomology ring $H^{\bullet}(F_{V})$.
Proof.
It follows from (35) that to prove Proposition 4.16, it is enough to show that for any permutation $\sigma\in\Sigma_{k}$ and any partition $D=A\sqcup B$ of $\{1,2,\ldots,k\}$, such that $A\neq\emptyset$, the polynomial $\phi_{D}(x_{1},\ldots,x_{k})$ defined in (41) restricts to zero in the cohomology ring $H^{\bullet}(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$.
Recall from Proposition 4.5 that $H^{\bullet}(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$ is isomorphic to the polynomial ring $\mathbb{Q}[x^{\sigma}_{1},\ldots,x^{\sigma}_{k}]$ quotient by the ideal $I_{\sigma}$ generated by
$\phi_{1}(x^{\sigma}_{\sigma(1)},\ldots,x^{\sigma}_{\sigma(k)})=(x^{\sigma}_{\sigma(1)})^{r_{\sigma(1)}}\text{\,\,\, and\,\,\,}\\ \phi_{i}(x^{\sigma}_{\sigma(1)},\ldots,x^{\sigma}_{\sigma(k)})=(x^{\sigma}_{\sigma(i)})^{r_{\sigma(i)}}\prod_{j=1}^{i-1}(x^{\sigma}_{\sigma(i)}-x^{\sigma}_{\sigma(j)})^{r_{\sigma(i)}r_{\sigma(j)}(g-1)}\text{\,\, for\,\,\,}2\leq i\leq k.$
Assume that $\phi_{D}(x^{\sigma}_{1},\ldots,x^{\sigma}_{k})\in\mathbb{Q}[x^{\sigma}_{1},\ldots,x^{\sigma}_{k}]$ is not zero for some $D=A\sqcup B$, then in particular, it is not divisible by $\phi_{1}(x^{\sigma}_{\sigma(1)},\ldots,x^{\sigma}_{\sigma(k)})$, hence $\sigma(1)\in B$. Since the polynomial $\phi_{D}(x^{\sigma}_{1},\ldots,x^{\sigma}_{k})$ is not divisible by $\phi_{2}(x^{\sigma}_{\sigma(1)},\ldots,x^{\sigma}_{\sigma(k)})$, we obtain that $\sigma(2)$ is also an element of $B$. Repeating the argument, we arrive at $B=\{\sigma(1),\sigma(2),\ldots,\sigma(k)\}$, and thus $A=\emptyset$. ∎
Let $I\subset\mathbb{Q}[x_{1},\ldots,x_{k}]$ be the ideal generated by the polynomials in (41) and let
$R=\mathbb{Q}[x_{1},\ldots,x_{k}]/I.$
It follows from Proposition 4.16 that sending the variable $x_{i}\in R$ to the class $x_{i}\in H^{\bullet}(F_{V})$ defined in Lemma 4.15 results in a ring morphism
$\phi:R\to H^{\bullet}(F_{V}).$
Our goal is to show that $\phi$ is an isomorphism. In order to show this, we use the following statement, which we will prove in §6.4.
Proposition 4.17.
Let $x_{i}\in R$ has degree 2 for any $1\leq i\leq k$, then the ring $R$ is a graded ring, whose Hilbert function coincides with the Poincaré polynomial of the fiber $F_{V}$.
Then, to prove that $\phi$ is an isomorphism, it is sufficient to prove injectivity.
Lemma 4.18.
The map $\phi:R\to H^{\bullet}(F_{V})$ is injective.
Proof.
To prove the statement, we need to show that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset I$. Let $d(\rho)=\sum_{1\leq i<j\leq k}r_{i}r_{j}(g-1)$; we start with proving that the surjective morphism of graded rings $\eta$:
[…]
induces an isomorphism between the homogeneous degree $2d(\rho)+2r-2k$ elements. It follows from Proposition 4.17 and the recursive formula (36), that the ring $R$ has $k!$ elements of top degree $2d(\rho)+2r-2k$. Since the top-dimensional cohomology of an equidimensional union of projective varieties equals the direct sum of the pieces, we have (see Propositions 4.1 and 4.10(i))
(42) $H^{2d(\rho)+2r-2k}(F_{V})\cong\bigoplus_{\sigma\in\Sigma_{k}}H^{2d(\rho)+2r-2k}(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}).$
Then using Lemma 4.15 and the integral formula (34), we obtain that for every $\sigma\in\Sigma_{k}$ the class
$\phi_{\sigma}(x_{1},\ldots,x_{k})\overset{\mathrm{def}}{=}\prod_{i=1}^{k}\Big{(}x_{\sigma(i)}^{r_{\sigma(i)}-1}\prod_{j>i}(x_{\sigma(j)}-x_{\sigma(i)})^{r_{\sigma(j)}r_{\sigma(i)}(g-1)}\Big{)}{\in H^{2d(\rho)+2r-2k}(F_{V})}$
maps to a non-zero generator of
$H^{2d(\rho)+2r-2k}(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$
under the $\sigma$-th coordinate projection in the direct sum (42). Since the morphism $\theta$ is injective (see the proof of Lemma 4.11), and the $k!$-dimensional space
$\oplus_{\sigma\in\Sigma_{k}}\mathbb{Q}\phi_{\sigma}(x_{1},\ldots,x_{k})\subset H^{2d(\rho)+2r-2k}(F_{V})$
is in the image of $\theta$, we conclude that the quotient ring $\mathbb{Q}[x_{1},\ldots,x_{k}]/\cap_{\sigma\in\Sigma_{k}}I_{\sigma}$ has at least $k!$ linearly independent elements of degree $2d(\rho)+2r-2k$.
Next, we repeat the argument from [17, proof of Theorem 7.1]: we note that the ring $R$ can be interpreted as the cohomology ring of the Lawrence toric variety of dimension $d(\rho)+r-k$ (cf. [31, Theorem 6.3] and [17, Theorem 8.3]), and thus the socle of $R$ consists precisely of the elements of top degree $2d(\rho)+2r-2k$. We assume that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\not\subset I$ and pick a non-zero element $h\in\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset R$ of maximal degree. Our goal is to show that $h$ is zero in $R$, and thus $h\in I$.
We note that since $h\in\cap_{\sigma\in\Sigma_{k}}I_{\sigma},$ we have also $f\cdot h\in\cap_{\sigma\in\Sigma_{k}}I_{\sigma}$ for any element $f\in R$. Clearly, for any non-constant $f\in R$ we have $\deg(f\cdot h)>\deg(h)$. By the maximality hypothesis in the choice of $h$, we obtain that $f\cdot h=0$ in $R$ and thus $f\cdot h$ belongs to $I$ for any non-constant $f\in R$. Hence $h$ is contained in the socle of $R$, and thus it has degree $2d(\rho)+2r-2k$. Finally, noting that $\eta(h)=0$, we conclude that $h$ is a zero element in $R$; this completes the proof. ∎
We end this section with a statement that we will need later, in §7.4. This formula uses Lemma 4.15, and is a result of a simple calculation, analogous to the one in the proof of Proposition 4.5.
Proposition 4.19.
Let $\sigma,\tau\in\Sigma_{k}$ and let $\mathcal{T}_{\sigma},\mathcal{T}_{\tau}$ be two towers of projective bundles defined in Proposition 4.13. The Chern character of the tangent bundle of $\mathcal{T}_{\tau}\cap\mathcal{T}_{\sigma}$ is
$ch(T(\mathcal{T}_{\tau}\cap\mathcal{T}_{\sigma}))=\sum_{\begin{subarray}{c}\tau^{-1}(i)<\tau^{-1}(j)\\ \sigma^{-1}(i)<\sigma^{-1}(j)\end{subarray}}r_{i}r_{j}(g-1)\exp(x_{j}-x_{i})+\sum_{1\leq i\leq k}r_{i}\exp(x_{i})-k.$
5. The no-support theorem for non-abelian strata
Finally, we turn to the general case of the fibers. Recall from §3.1 that a complete stratification $\mathcal{M}_{0}(r)=\sqcup_{\overline{\rho}}M_{\overline{\rho}}$ is parametrized by multi-partitions:
$\overline{\rho}=[(\bar{r}_{1},m_{1}),(\bar{r}_{2},m_{2}),\dots,(\bar{r}_{s},m_{s})]\vdash r;\;\;M_{\overline{\rho}}=\{\oplus_{i=1}^{s}V_{i}^{\oplus m_{i}}|\,V_{i}\in\mathcal{M}_{0}(r_{i})\},$
and we called a stratum $M_{\overline{\rho}}$ non-abelian if at least one multiplicity $m_{i}$ in $\overline{\rho}$ is greater than 1. The main goal of this section is to prove the following theorem.
Theorem 5.1.
A non-abelian stratum $M_{\overline{\rho}}$ carries no support of $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}$, thus the corresponding local system $\mathcal{L}_{\overline{\rho}}$ vanishes.
[…]
$L_{\overline{\rho}}=\mathrm{ker}\left(H^{\bullet}(F_{\overline{V}})\overset{j^{*}}{\longrightarrow}H^{\bullet}(\pi^{-1}(\mathcal{C}^{\circ}))\right).$
To prove the Theorem, we need to show that the map $j^{*}$ is injective.
In fact, we will prove a stronger statement. Recall from §3.1 that to the multi-partition $\overline{\rho}$ corresponds a true partition $\rho$ with $k=m_{1}+\dots+m_{s}$ elements. Now let $V\in M_{\rho}^{\mathrm{ab}}\cap\mathcal{C}^{\circ}$; we will show that the natural restriction map
$g:H^{\bullet}(F_{\overline{V}})\longrightarrow H^{\bullet}(F_{V})$
is injective, which clearly implies the injectivity of $j^{*}$.
To the multi-partition ${\overline{\rho}}$, we associate the subgroup $\Sigma[\overline{\rho}]\simeq\Sigma_{m_{1}}\times\dots\times\Sigma_{m_{s}}\subset\Sigma_{k}$, which exchanges the elements of ${\underline{k}}=\{1,2,\dots k\}$. The group $\Sigma[\overline{\rho}]$ acts naturally on the cohomology of an abelian fiber $F_{V}$, and, as we will show in Proposition 5.2, the image of the restriction map $g$ is contained in $H^{\bullet}(F_{V})^{\Sigma[\overline{\rho}]}$.
5.1. Injectivity and the cohomological map
The final statement is as follows.
Proposition 5.2.
Let $\overline{\rho}$ be a multi-partition and $\rho$ the associated partition, $\overline{V}\in M_{\overline{\rho}}$, and let $V\in M_{\rho}^{\mathrm{ab}}\cap U$ be as above. The map
(43) $g:H^{\bullet}(F_{\overline{V}})\longrightarrow H^{\bullet}(F_{V})^{\Sigma[\overline{\rho}]}$
is an isomorphism.
Proof.
To simplify our notation, we will describe in detail the case of $\overline{\rho}=[(s,k)]$, that is, the stratum $M_{\overline{\rho}}$ of bundles $\overline{V}=W^{\oplus k}$, where $W$ is a stable bundle of rank $s$ and degree 0. The general case is obtained by tensoring copies of this basic example (see Remark 5.12).
We thus have $sk=r$, and $\rho=[s,s,\dots,s]$. To avoid double indices, we introduce the notation $lW=W^{\oplus l}$. We note that the stratum $M_{\rho}^{\mathrm{ab}}$ may be obtained as a quotient of the space of sequences of stable bundles $\widetilde{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},\dots,V_{k})|\,V_{1}\oplus...\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}\}$ by the symmetric group $\Sigma_{k}$. Then the cohomologies of the abelian fibers $F_{\oplus V_{i}}$ form a trivial bundle over $\widetilde{M}_{\rho}^{\mathrm{ab}}$, to which the action of $\Sigma_{k}$ lifts as the permutation action on the variables $x_{j}$, $j=1,\dots,k$. It follows then that the image of the composed map (43) is contained in the invariant part of the cohomology of $H^{\bullet}(F_{V})$; we will show now that the map
(44) $g:H^{\bullet}(F_{kW})\longrightarrow H^{\bullet}(F_{V})^{\Sigma_{k}}$
is an isomorphism. Using induction, we assume that the statement holds for $k$ replaced by $k-1$, and our goal is to prove it for $k$. Just as in the abelian case, we begin with establishing the purity of the cohomology of the fiber.
Lemma 5.3.
The cohomology of $F_{kW}$ is pure.
Proof.
The argument is similar to the one given in the abelian case in the proof of Lemma 4.11. Our starting point is (23), which expresses the cohomology of $F_{kW}$ as a direct sum of contributitions from lower strata. Using an inductive hypothesis, we can assume that the lower non-abelian strata carry no support, and thus we can ignore them from the point of view of the Decomposition Theorem. This, up to taking products of our example, reduces the question of purity to the study of the contribution of the stratum $M_{\rho}^{\mathrm{ab}}$ with $\rho=[s,s,\dots,s]$. In this case, the monodromy group is $\Sigma_{k}$, and the variation of the Hodge structure of $H^{\bullet}(F_{V})$ as $V$ varies is trivial, since $H^{\bullet}(F_{V})$ has cohomology of Hodge-Tate type only (see (35)). Now, again we appeal to Proposition 3.14, and Corollary 3.17 to conclude that the cohomology of $F_{kW}$ is pure. ∎
A key object in our argument is the moduli space of stable parabolic extensions
^=$\hat{\mathcal{E}}=\{V^{\prime\prime}\to V\to V^{\prime}|\,V^{\prime\prime}\in\mathcal{M}_{0}(s),\,V^{\prime}\in\mathcal{P}_{0}((k-1)s),\,V\in\mathcal{P}_{0}(ks)\},$
which has two structure maps
(45) $\mathcal{P}_{0}((k-1)s)\overset{f^{\prime}}{\longleftarrow}\hat{\mathcal{E}}\overset{h^{\prime}}{\longrightarrow}\mathcal{P}_{0}(ks),$
where the map $f^{\prime}$ is a smooth fibration. In fact, we will be mainly interested by the subset ^$\hat{\mathcal{E}}_{\rho}$ of $\hat{\mathcal{E}}$ lying over the closure of $M_{\rho}^{\mathrm{ab}}\subset\mathcal{M}_{0}(ks)$.
We introduce the notation
$\widehat{\pi}:\hat{\mathcal{E}}\to\mathcal{M}_{0}(ks)$
for the composed map $\widehat{\pi}=\pi\circ h^{\prime}$. It follows from the proof of Proposition 4.1 that for an abelian $V=V_{1}\oplus\dots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ we have
(46) $\widehat{\pi}^{-1}(V)=\bigsqcup_{j=1}^{k}\mathcal{F}_{j},\;\;\;\mathcal{F}_{j}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\\ \sigma(k)=j\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})},\,j=1,\dots k.$
The purity of $H^{\bullet}(F_{V})$ (see Lemma 4.11) and that of the cohomology of the sets $\mathcal{F}_{j}$ (see (37)), which form a covering of $F_{V}$, imply that we have a $\Sigma_{k}$-invariant embedding $H^{\bullet}(F_{V})\hookrightarrow\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$, and this provides us with an embedding
$Q:H^{\bullet}(F_{V})^{\Sigma_{k}}\hookrightarrow\left[\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})\right]^{\Sigma_{k}}.$
We proceed analogously in the nonabelian case: for $\overline{V}=W^{\oplus k}$, we denote
=$\mathcal{E}\overset{\mathrm{def}}{=}\widehat{\pi}^{-1}(\overline{V})=\{0\to W\to W^{\prime\prime}\to W^{\prime}\to 0\;|\;W^{\prime}\in\pi^{-1}(W^{\oplus(k-1)}),\;W^{\prime\prime}\in\mathcal{P}_{0}(sk)\};$
restricting the map $h^{\prime}:\widehat{\mathcal{E}}\to\mathcal{P}_{0}(ks)$ (see (45)) to $\mathcal{E}\subset\widehat{\mathcal{E}}$, we obtain a birational morphism of projective varieties $h:\mathcal{E}\longrightarrow F_{kW}$.
Lemma 5.4.
The map $h^{*}:H^{\bullet}(F_{kW})\rightarrow H^{\bullet}(\mathcal{E})$ is injective.
Proof.
By Lemma 5.3, $H^{\bullet}(F_{kW})$ has pure Hodge structure, and so does $H^{\bullet}(\mathcal{E})$, since $\mathcal{E}$ is a bundle of projective spaces over $F_{(k-1)W}$. Now the injectivity follows from Proposition 3.13. ∎
Finally, we note that the cohomology $H^{\bullet}(\widehat{\pi}^{-1}(V))$ of the fibers (46) over $M_{\rho}^{\mathrm{ab}}$ may be obtained as a descent to $M_{\rho}^{\mathrm{ab}}$ of a trivial bundle $\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ on $\widetilde{M}_{\rho}^{\mathrm{ab}}$ endowed with a $\Sigma_{k}$-action. This results in a natural map
(47) $G:H^{\bullet}(\mathcal{E})\longrightarrow\left[\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})\right]^{\Sigma_{k}},$
which fits into the commutative diagram
[…]
Lemma 5.5.
The map $G$ from (47) is an isomorphism.
Proof.
Using descent with respect to the subgroup $\Sigma_{k-1}\subset\Sigma_{k}$, we can conclude that the projection of the map $H^{\bullet}(\mathcal{E})\longrightarrow\bigoplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ to $H^{\bullet}(\mathcal{F}_{1})$ has its image in $H^{\bullet}(\mathcal{F}_{1})^{\Sigma_{k-1}}$, and recall that by our inductive hypothesis
$H^{\bullet}(F_{(k-1)W})\simeq H^{\bullet}(F_{V/V_{1}})^{\Sigma_{k-1}}.$
Now, without loss of generality, we can assume that $V_{1}=W$, and then clearly $H^{\bullet}(\mathcal{E})\to H^{\bullet}(\mathcal{F}_{1})^{\Sigma_{k-1}}$ is an isomorphism, since $\mathcal{E}$ and $\mathcal{F}_{1}$ are projective bundles over $F_{(k-1)W}$ and $F_{V/V_{1}}$, respectively, which can be deformed into each other in a family.
Now the statement follows from the fact that the $\Sigma_{k}$-representation $\bigoplus_{i=j}^{k}H^{\bullet}(\mathcal{F}_{j})$ is induced from the representation $H^{\bullet}(\mathcal{F}_{1})$ of the subgroup $\Sigma_{k-1}\cong\mathrm{Stab}_{\Sigma_{k}}(1)$, and thus, by Frobenius reciprocity, we have $H^{\bullet}(\mathcal{F}_{1})^{\Sigma_{k-1}}\simeq\left[\bigoplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})\right]^{\Sigma_{k}}.$
Clearly, Lemma 5.5, and the commutative diagram (48) imply that the map (43) is injective. This already proves Theorem 5.1. For completeness, however, we will show below that this map is an isomorphism. ∎
5.2. The E-polynomial of non-abelian fibers
To show that the injection (43) is an isomorphism, it is enough to prove the following.
Proposition 5.6.
In the notation introduced above, we have
(49) $E_{t}(F_{\overline{V}})=\mathrm{Hilb}_{t}(H^{\bullet}(F_{V})^{\Sigma_{k}}).$
5.2.1. The right hand side
In preparation to the calculation of the cohomology of the non-abelian fibers, we refine and reformulate our motivic argument from the proof of Theorem 4.12 into a generalized Mayer-Vietoris argument.
Recall that we denote by $\mathcal{F}_{i}$ the union of projective towers
$\mathcal{F}_{i}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\\ \sigma(k)=i\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})},\,i=1,\ldots,k,$
and for a subset $J\subset\underline{k}$, we set
$\mathcal{F}_{J}=\bigcap_{i\in J}\mathcal{F}_{i}.$
Lemma 5.7.
For $i=1,\dots,k$, the inclusion $\mathcal{F}_{i}\hookrightarrow F_{V}$ induces a surjection in cohomology: $H^{\bullet}(F_{V})\longrightarrow H^{\bullet}(\mathcal{F}_{i})$, and, more generally, for $j\in J$, we have a surjection $H^{\bullet}(\mathcal{F}_{J\setminus\{j\}})\longrightarrow H^{\bullet}(\mathcal{F}_{J})$.
Proof.
Without loss of generality, we assume that $i=k$. Recall that in the proof of Theorem 4.12 we identified $\mathcal{F}_{k}$ with the projective bundle over the ”smaller” fiber of $\pi$ (see (37)). Thus using induction we can conclude that $H^{\bullet}(\mathcal{F}_{k})$ is generated by the $k$ classes $x_{1},\dots,x_{k}\in H^{2}(\mathcal{F}_{k})$. It follows from Lemma 4.15 that these classes are well-defined in the cohomology ring $H^{\bullet}(F_{V})$, hence the restriction map $H^{\bullet}(F_{V})\rightarrow H^{\bullet}(\mathcal{F}_{k})$ is surjective. ∎
Let $e_{1},\dots,e_{k}$ be the standard basis of $\mathbb{C}^{k}$; we associate $e_{j}$ to the component $V_{j}$ of $V=V_{1}\oplus\ldots\oplus V_{k}$. Then the set
$\{e_{J}=e_{j_{1}}\wedge\dots\wedge e_{j_{l}}|\,J=\{j_{1}<\dots<j_{l}\}\subset\underline{k}\}$
forms a basis of the exterior algebra $\wedge^{\bullet}\mathbb{C}^{k}$. For every $j\in\underline{k}$ we introduce the operators:
  • $\iota_{j}:\wedge^{m}\mathbb{C}^{k}\to\wedge^{m+1}\mathbb{C}^{k}$, (left) wedge product with $e_{j}$, and
  • with some abuse of notation, for any $J\subset\underline{k}$, we denote by $\kappa_{j}:H^{\bullet}(\mathcal{F}_{J\setminus\{j\}})\longrightarrow H^{\bullet}(\mathcal{F}_{J})$ the surjection from Lemma 5.7 if $j\in J$, and zero otherwise.
Now consider the operator
$Q=\sum_{j=1}^{k}\iota_{j}\otimes\kappa_{j}$
acting on the direct sum of vector spaces $H^{\bullet}_{J}=e_{J}\otimes H^{\bullet}(\mathcal{F}_{J})$.
Remark 5.8.
Note that $\mathbb{C}\cdot e_{J}$ is a 1-dimensional sign representation of the permutation of the elements of $J$.
The sets $\mathcal{F}_{i}$, $i=1,\dots,k$ form a cover of the fiber $F_{V}$, and the generalized Mayer-Vietoris sequence corresponding to this cover has the following form.
[…]
Proposition 5.9.
The generalized Mayer-Vietoris sequence (50) is exact.
Proof.
Indeed, recall that the spaces $\mathcal{F}_{J}$ are bundles of products of projective spaces over spaces of the form $F_{V^{\prime}}$. Since these latter spaces have cohomology with pure Hodge structure (see Lemma 4.11), so do the spaces $\mathcal{F}_{J}$. This immediately implies the exactness of (50). ∎
A key observation is that our operator $Q=\sum_{j=1}^{k}\iota_{j}\otimes\kappa_{j}$ is invariant under the action of $\Sigma_{k}$, and hence the invariant part of the complex (50) is an exact subcomplex of (50), whose first term is $H^{\bullet}(F_{V})^{\Sigma_{k}}$:
[…]
Lemma 5.10.
Let $l=1,\ldots,k-1$, then
$\left[\bigoplus_{|J|=l}H^{\bullet}_{J}\right]^{\Sigma_{k}}=H^{\bullet}(F_{(k-l)W})\otimes\wedge^{l}H^{\bullet}(\mathbb{P}R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}W)),$
where we denote by $\mathcal{U}_{k-l}$ the normalized (cf. def. 3.9) universal bundle over $F_{(k-l)W}\times C$, by $\xi$ and $\zeta$ the projections along $C$ and $F_{(k-l)W}$, respectively, and treat $W$ as a parabolic bundle with trivial parabolic structure. For the last term in (51) we have $H_{\underline{k}}^{\Sigma_{k}}\cong\wedge^{k}H^{\bullet}(\prod_{i=1}^{k}\mathbb{P}(W_{p}))$.
Remark 5.11.
Note that $R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}W))$ is a vector bundle over $F_{(k-l)W}$ of rank $(k-l)(g-1)s^{2}+s$.
Proof.
First, we observe that $\oplus_{|J|=l}H^{\bullet}_{J}$, as a representation of $\Sigma_{k}$, is induced from the representation $H^{\bullet}_{\{1,2,\dots,l\}}$ of the stabilizer subgroup $\Sigma_{l}\times\Sigma_{k-l}$, and thus by Frobenius reciprocity, we have
$\left[\bigoplus_{|J|=l}H^{\bullet}_{J}\right]^{\Sigma_{k}}=\left[H^{\bullet}_{\{1,2,\dots,l\}}\right]^{\Sigma_{l}\times\Sigma_{k-l}}.$
The space $\mathcal{F}_{\{1,2,\dots,l\}}$ is a product of projective bundles
$\prod_{j=1}^{l}\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{W}^{V_{l+1}\oplus\dots\oplus V_{k}},\zeta^{*}V_{j}))$
over the base $F_{V_{l+1}\oplus\dots\oplus V_{k}}$ (cf. the proof of Theorem 4.12). After taking the invariants under the $\Sigma_{k-l}$-action on the cohomology rings, using the inductive hypothesis, we arrive at the tensor product
$\bigotimes_{j=1}^{l}H^{\bullet}(\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}V_{j})))\otimes H^{\bullet}(F_{(k-l)W}),$
on which $\Sigma_{l}$ acts by permuting the first $l$ tensor factors. These factors are naturally identified then, and thus, below in (52), we will replace them with the same vector space. Taking the invariants of this $\Sigma_{l}$-representation, tensored with the sign representation (cf. Remark 5.8), we obtain the wedge power
(52) $\wedge^{l}H^{\bullet}(\mathbb{P}R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}W))\otimes H^{\bullet}(F_{(k-l)W}).$
The last statement of Lemma 5.10 follows from the observation: $\mathcal{F}_{\underline{k}}\cong\prod_{i=1}^{k}\mathbb{P}(W_{p})$ (cf. the proof of Theorem 4.12). ∎
Comparing (51) and Lemma 5.10, we obtain
(53) $\mathrm{Hilb}_{t}(H^{\bullet}(F_{V})^{\Sigma_{k}})=\sum_{l=1}^{k}(-1)^{l+1}E_{t}(F_{(k-l)W})\cdot E_{t}(\wedge^{l}H_{\bullet}(\mathbb{P}^{s^{2}(k-l)(g-1)+s-1})).$
Next, we consider the moduli space of extensions of parabolic bundles: for $l=1,\ldots,k-1$ we define
$\mathcal{E}_{l}\overset{\mathrm{def}}{=}\{0\to W^{\oplus l}\to W^{\prime\prime}\to W^{\prime}\to 0\;|\;W^{\prime}\in\pi^{-1}(W^{\oplus(k-l)})\}$
which is a natural generalization of the space $\mathcal{E}=\mathcal{E}_{1}$ introduced in §5.1. We note that
$\mathrm{Aut}(W^{\oplus l})\cong GL_{l}\text{\, and \,}\mathrm{Gr}_{l}(N)\cong\mathrm{Hom}(\mathbb{C}^{l},\mathbb{C}^{N})\sslash GL_{l},$
where we denote by $\mathrm{Gr}_{l}(N)$ the space of $l$-dimensional subspaces in $\mathbb{C}^{N}$. Hence it follows from Remark 5.11 that we have the equality of weight polynomials
(54) $E_{t}(\mathcal{E}_{l})=E_{t}(F_{(k-l)W})\cdot E_{t}(\mathrm{Gr}_{l}(s^{2}(k-l)(g-1)+s))$
for $l=1,\dots,k-1$. Recall the definition of the Gauss binomial coefficients:
$\binom{m}{n}_{t}=\frac{(1-t^{m})(1-t^{m-1})\ldots(1-t^{m-n+1})}{(1-t)(1-t^{2})\ldots(1-t^{n})}\text{\; for\;}n=1,\ldots,m;\;\;\binom{m}{0}_{t}=1.$
Note that
$E_{t}(\wedge^{n}H^{\bullet}(\mathbb{P}^{m-1}))=t^{\binom{n}{2}}\binom{m}{n}_{t},$
which follows from the Cauchy binomial theorem
(55) $\prod_{k=0}^{m-1}(1+t^{k}z)=\sum_{n=0}^{m}t^{\binom{n}{2}}\binom{m}{n}_{t}z^{n}.$
Since $E_{t}(\mathrm{Gr}_{n}(N))=\binom{N}{n}_{t}$, we have
(56) $E_{t}(\wedge^{n}H^{\bullet}(\mathbb{P}^{N-1}))=t^{\binom{n}{2}}\cdot E_{t}(\mathrm{Gr}_{n}(N)).$
Putting (53), (54) and (56) together, we obtain
(57) $\mathrm{Hilb}_{t}(H^{\bullet}(F_{V})^{\Sigma_{k}})=\sum_{l=1}^{k}(-1)^{l+1}t^{\binom{l}{2}}E_{t}(\mathcal{E}_{l}).$
5.2.2. The left hand side
Now we will show that the same structure appears when calculating the cohomology of the non-abelian fibers. We start with a detailed description of the fiber $F_{\overline{V}}$ for $\overline{V}\cong W\oplus W$.
Example 1.
The fiber $F_{W\oplus W}$. In this case, the space
=$\mathcal{E}\overset{\mathrm{def}}{=}\{0\to W\to W^{\prime\prime}\to W^{\prime}\to 0|\;W^{\prime}\in\pi^{-1}(W),\;W^{\prime\prime}\in\mathcal{P}_{0}(2s)\}$
defined in §5.1 is a projective bundle with fibers $\mathbb{P}^{s^{2}(g-1)+s-1}$ over the projective space $\mathbb{P}(W_{p})\cong\mathbb{P}^{s-1}$, which corresponds to the choice of a line in $W_{p}$, and then the choice of an appropriate lift of this line to the fiber $W^{\prime\prime}_{p}$ of a vector bundle $W^{\prime\prime}$ obtained as an extension of $W$ by $W$ (see Proposition 4.1).
As opposed to the abelian case, some of parabolic extensions $(W^{\prime\prime},l)$ obtained in such way, are isomorphic, however. Indeed, consider the locus $\mathcal{G}\subset\mathcal{E}$ of those parabolic bundles $W^{\prime\prime}$ which, as vector bundles, are isomorphic to $W\oplus W$. The relevant parabolic structure on $W^{\prime\prime}\in\mathcal{G}$ is a choice of a line $l$ in $W_{p}$, and then an appropriate lift of this line to $W_{p}\oplus W_{p}$. Thus it is easy to see that the locus $\mathcal{G}\subset\mathcal{E}$ forms the space of (1,2)-flags of $W_{p}$.
However, a simple calculation shows that a parabolic structure on $W\oplus W$ is given by the choice of two non-colinear vectors in $W_{p}$, up to the action of $\mathrm{Aut}(W\oplus W)\cong GL_{2}$. Hence the space of non-isomorphic parabolic structures on $W\oplus W$ forms the Grassmannian $\mathrm{Gr}_{2}(W_{p})$.
The final picture of the fiber $F_{W\oplus W}$ is the following: over a base $\mathbb{P}^{s-1}$, there is a projective bundle with fibers $\mathbb{P}^{s^{2}(g-1)+s-1}$, which contains a subbundle $\mathcal{G}$, which forms a (1,2)-flag variety of the vector space $W_{p}$. The projection $\mathcal{G}\to\mathrm{Gr}_{2}(W_{p})$ is a $\mathbb{P}^{1}$-fibration, and to obtain the fiber $F_{W\oplus W}$, we need to collapse this fibration to the base.
Now, we generalize the example described above. We observe that the fiber $F_{kW}$ has a decomposition into locally closed subsets
(58) $F_{kW}=\mathcal{S}_{1}\sqcup\mathcal{S}_{2}\sqcup\dots\sqcup\mathcal{S}_{k},$
where $\mathcal{S}_{l}$ is the subset of those parabolic bundles which have $W^{\oplus l}$ as a subbundle, but not $W^{\oplus(l+1)}$. In particular, we note that $\mathcal{S}_{k}\cong{\mathrm{Gr}_{k}(W_{p})}$.
On the other hand, decomposing $\mathcal{E}_{l}$ according to which element of (58) $W^{\prime}$ belongs, we obtain
$\mathcal{E}_{l}=\bigsqcup_{m\geq l}\mathcal{S}_{m}\times\mathrm{Gr}_{l}(m),$
where $\mathrm{Gr}_{l}(m)$ represents a choice of $W^{\oplus l}\subset W^{\oplus m}$. Using the motivic property of the E-polynomial, Proposition 3.19, we obtain
$E_{t}(\mathcal{E}_{l})=\sum_{m\geq l}E_{t}(\mathcal{S}_{m})\cdot E_{t}(\mathrm{Gr}_{l}(m)).$
Substituting this into (57), gives us
$\mathrm{Hilb}_{t}(H^{\bullet}(F_{V})^{\Sigma_{k}})=\sum_{l=1}^{k}(-1)^{l+1}\sum_{m\geq l}E_{t}(\mathcal{S}_{m})\cdot t^{\binom{l}{2}}\cdot E_{t}(\mathrm{Gr}_{l}(m)).$
Finally, noting that setting $z=-1$ in (55) reduces the formula to
$1=\sum_{n=1}^{m}(-1)^{n+1}t^{\binom{n}{2}}\binom{m}{n}_{t},$
we obtain:
$\mathrm{Hilb}_{t}(H^{\bullet}(F_{V})^{\Sigma_{k}})=\sum_{m=1}^{k}E_{t}(\mathcal{S}_{m})\sum_{l=1}^{m}(-1)^{l+1}\cdot t^{\binom{l}{2}}\cdot E_{t}(\mathrm{Gr}_{l}(m))=\\ \sum_{m=1}^{k}E_{t}(\mathcal{S}_{m})\sum_{l=1}^{m}(-1)^{l+1}\cdot t^{\binom{l}{2}}\cdot\binom{m}{l}_{t}=\sum_{m=1}^{k}E_{t}(\mathcal{S}_{m}),$
which equals $E_{t}(F_{kW})$, since we have the decomposition (58).
Remark 5.12.
The case of general non-abelian fibers corresponds to tensoring the relevant complexes described above, and thus follows from the proof presented for the case $W^{\oplus k}$.
6. Combinatorics of subtraction
Moving to the next step of our plan (see §3.6), in this section we begin the study of the stalk of the local system over the abelian stratum $M_{\rho}^{\mathrm{ab}}$ contributing to the Decomposition Theorem.
Recall our simplified notation from §3.6. Fix a point $V\in M_{\rho}^{\mathrm{ab}}$; we write
  • $F_{\rho}=\pi^{-1}(V)$ for the fiber of $\pi$, and
  • $L_{\rho}=(\mathcal{L}_{\rho})_{V}$ for the stalk of the local system supported at $V$; we thus have $L_{\rho}\subset H^{\bullet}(F_{\rho})$.
Our goal in this section is to calculate the Hilbert function of the graded vector space $L_{\rho}\subset H^{\bullet}(F_{\rho})$, using equality (1.7.B) of the Decomposition Theorem. Applied to our situation the formula has the form
(59) $H^{\bullet}(F_{\rho})=L_{\rho}\oplus\bigoplus_{\mu>\rho}IH^{\bullet}(\mathcal{C}_{\mu\rho},\mathcal{L}_{\mu}),$
where $\mathcal{C}_{\mu\rho}$ denotes the normal slice to $M_{\rho}$ in $M_{\mu}$ at the point $V$.
In particular, the section is devoted to the proof of the following result.
Theorem 6.1.
For $\rho=[r_{1},\ldots,r_{k}]\vdash r$, the graded vector space $L_{\rho}\subset H^{\bullet}(F_{\rho})$ is the stalk of the local system $\mathcal{L}_{\rho}$ supported on the stratum $M_{\rho}^{\mathrm{ab}}$. Then the Hilbert function $\mathrm{Hilb}_{t}(L_{\rho})=\sum_{i}\dim L^{i}_{\rho}\cdot t^{i}$ is given by
(60) $\mathrm{Hilb}_{t}(L_{\rho})=t^{2d(\rho)}\prod_{i=1}^{k}p(r_{i};t^{2}),$
where $d(\rho)=\sum_{i<j}r_{i}r_{j}(g-1)$, and $p(n;t)=(t^{n}-1)/(t-1)$.
6.1. The subtraction algorithm
As a first step, we derive a recursion (the subtraction algorithm) for $\mathrm{Hilb}_{t}(L_{\rho})$ based on (59). To describe the structure of the cone $\mathcal{C}_{\mu\rho}$, we need to introduce a combinatorial device.
Definition 6.2.
A decomposition $\underline{\lambda}=\{\underline{\lambda}_{1},\ldots,\underline{\lambda}_{n}\}$ of $\underline{k}=\{1,\ldots,k\}$ is a representation of $\underline{k}$ as a disjoint union of subsets; we will use the notation $\underline{\lambda}\vdash\underline{k}$ in this case.
Given $\underline{\lambda}=\{\underline{\lambda}_{1},\ldots,\underline{\lambda}_{n}\}\vdash\underline{k}$ and a partition $\rho=[r_{1},\ldots,r_{k}]\vdash r$, we can compose $\underline{\lambda}$ and $\rho$ to obtain a new partition $\underline{\lambda}\circ\rho=\mu=[m_{1},\ldots,m_{n}]\vdash r$ such that $m_{j}=\sum_{i\in\underline{\lambda}_{j}}r_{i}$ for $j=1,\ldots,n$; we will denote this partition of $m_{j}$ by $\rho|\underline{\lambda}_{j}$.
Given a polystable bundle $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ and $\underline{\lambda}\vdash\underline{k}$, then we can define the bundles
$V[\underline{\lambda}_{j}]=\bigoplus_{i\in\underline{\lambda}_{j}}V_{i}\in\mathcal{M}_{0}(m_{j}),\quad j=1,\ldots,n,$
where $\mu=[m_{1},\ldots,m_{n}]=\underline{\lambda}\circ\rho$. Note that by construction $\rho$ is a refinement of $\mu$. Also, it is easy to see that any partition of $r$ having $\rho$ as a refinement can be written in this way.
Now recall the notation $\mathcal{C}_{V}$ for the normal slice at $V\in\mathcal{M}_{0}(r)$ of the stratum containing $V$ (cf. Proposition 3.7). For $V$ abelian, up to isomorphism, this normal slice only depends on the stratum $M_{\rho}$ to which $V$ belongs. If this causes no confusion, we will simply write $\mathcal{C}_{\rho}$ instead of $\mathcal{C}_{V}$.
The following technical statements are straightforward and their proof will be omitted.
Lemma 6.3.
Suppose $\mu=\underline{\lambda}\circ\rho$, and $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$. Then $M_{\rho}\subset\overline{M}_{\mu}$ and
  1. (1)
    $\mathcal{C}_{\mu\rho}\cong\bigcup_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}\prod_{j=1}^{n}\mathcal{C}_{\rho|\underline{\lambda}_{j}},\text{\; and}$
  2. (2)
    the normalization of $\mathcal{C}_{\mu\rho}$ is the corresponding disjoint union:
    $\tilde{\mathcal{C}}_{\mu\rho}\cong\bigsqcup_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}\prod_{j=1}^{n}\mathcal{C}_{\rho|\underline{\lambda}_{j}}.$
  3. (3)
    The local system $\mathcal{L}_{\mu}$ is trivial on $M_{\mu}$ in a neighborhood of $V$.
These statements imply that
$IH^{\bullet}(\mathcal{C}_{\mu\rho},\mathcal{L}_{\mu})=\sum_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}L_{\mu}\otimes\bigotimes_{j=1}^{n}IH^{\bullet}(\mathcal{C}_{\rho|\underline{\lambda}_{j}}).$
Now we can take the Hilbert function of the two sides of (59), and arrive at the following identity
$P_{t}(F_{\rho})=\mathrm{Hilb}_{t}\left(L_{\rho}\right)+\sum_{\mu>\rho}\sum_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}\mathrm{Hilb}_{t}\left(L_{\mu}\right)\cdot\prod_{j=1}^{n}IP_{t}(\mathcal{C}_{\rho|\underline{\lambda}_{j}}).$
Finally, we can rewrite this formula slightly differently, and arrive at our subtraction formula.
Proposition 6.4.
Given an abelian point $V\in M_{\rho}^{\mathrm{ab}}$, the Hilbert function of the fiber of the local system at $V$ may be expressed via the following recursive formula:
(61) $\mathrm{Hilb}_{t}\left(L_{\rho}\right)=P_{t}(F_{\rho})-\sum_{\underline{\lambda}\vdash\underline{k}}\mathrm{Hilb}_{t}\left(L_{\underline{\lambda}\circ\rho}\right)\cdot\prod_{j=1}^{n}IP_{t}(\mathcal{C}_{\rho|\underline{\lambda}_{j}}).$
This recursion requires a lot of data: we need to know both $P_{t}(F_{\rho})$ and $IP_{t}(\mathcal{C}_{\rho})$ for all partitions $\rho\vdash r$ in order to calculate $\mathrm{Hilb}_{t}\left(L_{\rho}\right)$. The cohomology of $F_{\rho}$ was described in §4, while the intersection cohomology of the affine toric variety $\mathcal{C}_{\rho}$ was studied in detail in [31]. To prove Theorem 6.1, we will need to reconcile these different descriptions, and we will do this using the combinatorics of certain oriented graphs. In what follows, we will liberally draw on the ideas of [31].
6.2. Acyclic subgraphs and intersection cohomology
We will now describe the main result of [31]: a combinatorial model of $IP_{t}(\mathcal{C}_{\rho})$, $\rho\vdash r$.
Definition 6.5.
Let $\Gamma$ be an oriented graph on a finite vertex set $V(\Gamma)$, and with possibly multiple edges, but no loops; we denote the set of edges of $\Gamma$ by $E(\Gamma)$, and their number $|E(\Gamma)|$ by $e(\Gamma)$.
We will say $\Gamma$ is rooted in $v\in V(\Gamma)$ if from any vertex, there is an oriented path to $v$; we will say $\Gamma$ is acyclic if it has no oriented paths starting and ending at the same vertex. We introduce the notation
$\mathrm{Acyc}[\Gamma,v,n]=\{G\subset\Gamma|\ V(G)=V(\Gamma),G\text{ acyclic, rooted in }v,e(G)=n\}$
for the set of acyclic subgraphs of $G$ rooted in $v$, and having $n$ edges.
Finally, by generating function of a set of graphs $\mathcal{A}$ we will always mean counting the graphs weighted by the number of their edges: usually, of the form $t^{N}\cdot\sum_{\Gamma\in\mathcal{A}}t^{e(\Gamma)}$ for some integer $N$.
Now we introduce two graphs associated to partitions of $r$:
  • Given a partition $\rho=[r_{1},..,r_{k}]\vdash r$, we denote by $\mathbb{F}_{\rho}$ the directed graph with vertex set $\{0,1,2,\ldots,k\}$, with no out-edges from the zero vertex, with number of oriented edges from $i$ to 00 equal to $r_{i}$, and with number of oriented edges from $i$ to $j$ equal to $r_{i}r_{j}(g-1)$ for $0<i\neq j\leq k$.
  • We denote by $\mathbb{F}^{\circ}_{\rho}\subset\mathbb{F}_{\rho}$ the subgraph obtained from $\mathbb{F}_{\rho}$ by removing the vertex labelled by 00 and all edges attached to this vertex.
The following theorem of [31] gives a model of the coefficients of the shifted ($t\to t+1$) intersection Poincaré polynomial of $\mathcal{C}_{\rho}$.
Theorem 6.6.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, and denote by $\mathcal{C}_{\rho}$ the affine toric variety defined in (18). Introduce the notation $g(\rho;t^{2})=IP_{t}(\mathcal{C}_{\rho},t)$. Then, for any vertex $q\in\underline{k}$ of the graph $\mathbb{F}^{\circ}_{\rho}$, we have
(62) $g({\rho};t+1)=\sum_{i=1}^{d(\rho)-k+1}\left|\mathrm{Acyc}[\mathbb{F}^{\circ}_{\rho},q,k-1+i]\right|\cdot t^{i},$
where $d(\rho)=\sum_{i<j}(g-1)r_{i}r_{j}$.
6.3. Poincaré polynomials of the fibers in terms of graphs
Now we give a similar graph-theoretic interpretation for the Poincaré polynomials of the abelian fibers of $\pi$.
Theorem 6.7.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, $V\in M_{\rho}^{\mathrm{ab}}$ and recall the notation $f(\rho;t^{2})$ for the Poincaré polynomial of the fiber $\pi^{-1}(V)$. Then
(63) $f(\rho;t+1)=\sum_{i=1}^{d(\rho)+r-k}\left|\mathrm{Acyc}[\mathbb{F}_{\rho},0,k+i]\right|\cdot t^{i}.$
Proof.
We will prove the theorem using induction on $k$. A simple calculation shows that the statement is true for $k=1$.
Performing a change of variables $t\to t+1$ in the recursion from Theorem 4.12, we obtain that
$f(\rho;t+1)=\sum_{J\subset\{1,\ldots,k\}}(-1)^{|J|-1}f(\rho|{\overline{J}};t+1)\prod_{j\in J}((t+1)^{r_{j}r_{\overline{J}}(g-1)+r_{j}}-1)/t,$
where the sum is taken over all non-empty subsets $J=\{j_{1},\ldots,j_{s}\}\subset\{1,2,\ldots,k\}$, $\bar{J}=\{1,2,\ldots,k\}\setminus J$, and $\rho|J=[r_{j_{1}},\ldots,r_{j_{s}}]\vdash r_{J}$, where $r_{J}=\sum_{j\in J}r_{j}$. We start the recursion with $f(\rho|{\emptyset};t)=1$, $r_{\emptyset}=0$.
Now fix a nonempty subset $J\subset\{1,2,\ldots,k\}$. Let $G$ be a connected acyclic subgraph of $\mathbb{F}_{\rho}$ with $k+1$ vertices, rooted at zero vertex, and such that the vertices $j\in J$ of $G$ have no in-edges. Observe that
  • the polynomial $((t+1)^{r_{j}r_{\overline{J}}(g-1)+r_{j}}-1)/t$ counts all out-edges of $G$ from the vertex $j\in J$;
  • the graph $G\setminus J$ obtained from $G$ by deleting all vertices labelled by $j\in J$ and all edges attached to these vertices is an acyclic subgraph of $\mathbb{F}_{\rho}$ rooted at zero vertex, with vertex set $\overline{J}\cup\{0\}$;
  • it follows from the induction hypothesis that the edges of $G\setminus J$ are counted by the polynomial $f(\rho|{\overline{J}};t+1)$.
Noting that any edge of $G$ is either an out-edge from some vertex $j\in J$, or is an edge of $G\setminus J$, and using induction on $k$, we arrive at the equality
(64) $f(\rho|{\overline{J}};t+1)\prod_{j\in J}((t+1)^{r_{j}r_{\overline{J}}(g-1)+r_{j}}-1)/t=\sum_{G}t^{e(G)-k},$
where the sum is taken over the set of connected acyclic subgraphs $G\subset\mathbb{F}_{\rho}$ with $k+1$ vertices, rooted at the zero vertex, and such that the vertices $j\in J$ of $G$ have no in-edges. Now the recursion (63) follows from the inclusion-exclusion principle. ∎
Now we are ready to prove the main combinatorial result of this section.
Theorem 6.8.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$. Recall the notation introduced above: for $n\in\mathbb{N}$ we set $p(n;t)=(t^{n}-1)/(t-1)$ and for $\underline{\lambda}=(\underline{\lambda}_{1},\dots,\underline{\lambda}_{n})\vdash\underline{k}$, we denote by $\underline{\lambda}\circ\rho$ the corresponding composed partition of $r$. We set $r_{\underline{\lambda}_{j}}=\sum_{i\in\underline{\lambda}_{j}}r_{i}$ and denote by $\rho|\underline{\lambda}_{j}$ the partition of $r_{\underline{\lambda}_{j}}$ induced by $\rho$. Finally, we set $d(\underline{\lambda}\circ\rho)=\sum_{i<j}r_{\underline{\lambda}_{i}}r_{\underline{\lambda}_{j}}(g-1)$. Then
(65) $f(\rho;t)=\sum_{\underline{\lambda}\vdash\underline{k}}t^{d(\underline{\lambda}\circ\rho)}\prod_{i}g(\rho|{\underline{\lambda}_{i}};t)p(r_{\underline{\lambda}_{i}};t),$
where the sum is taken over all partitions $\underline{\lambda}=(\underline{\lambda}_{1},\ldots,\underline{\lambda}_{n})$ of the set $\underline{k}=\{1,2,\ldots,k\}$.
Proof.
First, we perform a change of variable $t\mapsto t+1$ and rewrite equation (65) as
(66) $f(\rho;t+1)=\sum_{\underline{\lambda}\vdash\underline{k}}(t+1)^{d(\underline{\lambda}\circ\rho)}\prod_{i}g(\rho|{\underline{\lambda}_{i}};t+1)p(r_{\underline{\lambda}_{i}};t+1).$
Our task is to construct a bijection between the combinatorial model of the left hand side, given in Theorem 6.7, and the very similar model on the right hand side given in Theorem 6.6, and this requires some care.
Let us note that the combinatorial choices on the right hand side involve
  • a decomposition $\underline{\lambda}=\{\underline{\lambda}_{1},\ldots,\underline{\lambda}_{n}\}\vdash\underline{k}$,
  • rooted acyclic graphs $G_{m}$ with vertex sets $\underline{\lambda}_{m}$,
  • a set of edges pointing towards the vertex 0 (informally, we will call these red); once $\underline{\lambda}$ is fixed, the generating function for the number of these edges (at least 1, for each $\underline{\lambda}_{m}$) is precisely $\prod_{m}p(r_{\underline{\lambda}_{m}};t+1),$
  • a set of edges between the different $\underline{\lambda}_{j}$ (informally, we will call these blue); under the hypothesis that we permit edges pointing from $\underline{\lambda}_{i}$ to $\underline{\lambda}_{j}$ only if $i<j$, the generating function of the number of these edges is precisely $(t+1)^{d(\underline{\lambda}\circ\rho)}$.
In the construction of the bijection, we will take advantage of the freedom of choice of the root-vertex provided by Theorem 6.6.
We introduce some simple notation to help describe our construction: let $v,w$ be two vertices of an oriented graph $\Gamma$
  • we will write $(v\to w)$ if there is an oriented edge pointing from $v$ to $w$;
  • we will write $(v\rightsquigarrow w)$ if there is an oriented path from $v$ to $w$.
In particular, if $\Gamma$ is rooted in $w$, then we have: $\forall v\;(v\rightsquigarrow w)$. We will assume that $(v\rightsquigarrow v)$ is true, but $(v\to v)$ is not.
Finally, if $S\subset V(\Gamma)$, then $\Gamma|S$ is the oriented graph whose vertex-set is $S$, and for which $v,w\in S$, $(v\to w)$ in $\Gamma|S$ if and only if $(v\to w)$ in $\Gamma$.
[…]
Let $\Gamma\in\mathrm{Acyc}[\mathbb{F}_{\rho},0,n]$ (see Theorem 6.7). We will call a vertex $j$ red if $(j\to 0)$; the set of red vertices is thus $\mathrm{Red}(\Gamma)=\{j\in\underline{k}\;|\;(j\to 0)\}$.
Now let $j_{1}$ be the smallest element of $\mathrm{Red}(\Gamma)$, and define $\underline{\lambda}_{1}=\{j\in\underline{k}|\;(j\rightsquigarrow j_{1})\}$, and $G_{1}=\Gamma|\underline{\lambda}_{1}$.
Next we cut $G_{1}$ from $\Gamma$ and find the smallest red vertex $j_{2}$ in $\Gamma|\{\{0,1,2,\dots,k\}\setminus\underline{\lambda}_{1}\}$. Again, we define $\underline{\lambda}_{2}=\{j\in\underline{k}\;|\;j\notin\underline{\lambda}_{1},\,(j\rightsquigarrow j_{2})\}$, and $G_{2}=\Gamma|\underline{\lambda}_{2}$, and we continue this process until we exhaust $\underline{k}$. We note that by construction
  • each $G_{m}$ is acyclic graph with vertex set $\underline{\lambda}_{m}$,
  • $G_{m}$ is rooted in $j_{m}$,
  • in $\underline{\lambda}_{m}$ the smallest vertex connected to 0 is $j_{m}$, and
  • if $v\in\underline{\lambda}_{i},w\in\underline{\lambda}_{j}$ and $(v\to w)$, then $i<j$.
[…]
Clearly, these two maps are inverses of each other, and thus establish a bijection between the two sets of choices. This completes the proof. ∎
We are now ready to present the proof of Theorem 6.1: Substituting $g(\rho;t^{2})=IP_{t}(\mathcal{C}_{\rho},t)$ and $f(\rho;t^{2})=P_{t}(F_{\rho})$ (cf. Theorems 6.6 and 6.7) into (61), we obtain a recursion
$\mathrm{Hilb}_{t}\left(L_{\rho}\right)=f(\rho;t^{2})-\sum_{\underline{\lambda}\vdash\underline{k}}\mathrm{Hilb}_{t}\left(L_{\underline{\lambda}\circ\rho}\right)\cdot\prod_{j=1}^{n}g(\rho|\underline{\lambda}_{j};t^{2})$
for the functions $\mathrm{Hilb}_{t}\left(L_{\rho}\right)$, $\rho\vdash r$. The recursion starts with the trivial partition: $\rho=[r]$; in this case, $\mathrm{Hilb}_{t}\left(L_{[r]}\right)=P_{t}(F_{[r]})$ is the cohomology of the generic fiber of the map $\pi$, which is isomorphic to $\mathbb{P}^{r-1}$, and thus $\mathrm{Hilb}_{t}\left(L_{[r]}\right)=p(r;t^{2})$. Now we observe that this recursion is identical to (65), and has the same starting condition, and thus has to have the same solution: $\mathrm{Hilb}_{t}\left(L_{\rho}\right)=t^{2d(\rho)}\prod_{i=1}^{k}p(r_{i};t^{2}).$
6.4. Proof of Proposition 4.17
As an added benefit of our combinatorial models, we can now prove Proposition 4.17, which states that $f(\rho;t^{2})$, the Poincaré polynomial of $F_{\rho}$ coincides with the Hilbert function of a concrete ring, given by generators and relations.
First note that the polynomial $g(\rho;t)$ from (62) depends only on the graph $\mathbb{F}^{\circ}_{\rho}$, $g(\rho;t)=\hat{g}(\mathbb{F}^{\circ}_{\rho},t)$, and one can define $\hat{g}(\Gamma,t)$ for any other directed graph on a vertex set $V(\Gamma)$ and with $r_{ij}=r_{ji}$ edges pointing from vertex $i$ to $j$ and $j$ to $i$.
The key point is an observation of [31, Theorem 6.3]: the function $\hat{g}(\Gamma,t^{2})$ coincides with the Hilbert function of the quotient algebra
(67) $\mathbb{Q}[x_{i}-x_{j},0\leq i<j\leq k]/\big{\langle}\psi_{D}=\prod_{\begin{subarray}{c}i\in A,j\in B\end{subarray}}(x_{i}-x_{j})^{r_{ij}},\;D=A\sqcup B\big{\rangle},$
where $D=A\sqcup B$ runs over all nontrivial partitions of the vertex set $V(\Gamma)$ (cf. [17]).
Now we note that the polynomial $f(\rho;t)$ from (63) coincides with the polynomial $\hat{g}(\hat{\mathbb{F}}_{\rho},t)$, where $\hat{\mathbb{F}}_{\rho}$ is the same graph as $\mathbb{F}_{\rho}$, but with some useless edges added: we set $r_{0i}=r_{i0}=r_{i}$, for $i=1,\dots,k$ and declare the root to be the vertex 00. We have then
$f(\rho;t)=\hat{g}(\hat{\mathbb{F}}_{\rho},t),$
and thus $f(\rho;t^{2})$ equals the Hilbert function of (67) for the data $r_{ij}=r_{ji}=r_{i}r_{j}(g-1)$ for $0<i<j\leq k$, and $r_{0i}=r_{i0}=r_{i}$, $i=1,\dots,k$. Now we can complete the proof of Proposition 4.17 by noting that the change of variables $x_{i}-x_{0}\mapsto x_{i}$, $i=1,\ldots,k$ converts the algebra (67) to the algebra (41).
7. Calculation of the local systems
In this section, we will describe the local system $\mathcal{L}_{\rho}$ appearing in (24) (cf. (1.7.A)) by identifying its fiber $L_{\rho}$ (cf. (25)) as a subspace of $H^{\bullet}(F_{\rho})$, see (75) for the result.
As a first step, we recast this problem in the framework of the calculation of certain enumerative quantities, the so called refined intersection forms. Let us briefly recall what these are in a somewhat more general context.
7.1. Refined intersection forms
Assume that $Z$ is a projective, possibly singular subvariety in a smooth not necessarily compact variety $X$ of dimension $m\leq 2\dim Z$. Let $S,T\subset Z$ be cycles representing classes $[S]\in H_{n}(Z)$ and $[T]\in H_{2m-n}(Z)$, respectively.
Denote the class of $S$ in $H_{n}(X)$ by $[S]_{X}$, and the class of $T$ in $H_{2m-n}(X)$ by $[T]_{X}$. A key object is the map $H_{\bullet}({X})\overset{\Psi}{\longrightarrow}H_{\bullet}^{\mathrm{BM}}(X)$ from homology to locally finite (or Borel-Moore) homology, which is dual to the mapping of compactly supported cohomology classes to ordinary cohomology.
[…]
(68) $[S]\cdot_{X}[T]=\langle[S]_{X},\Psi([T]_{X})\rangle\in\mathbb{Q}.$
This pairing is called the refined intersection form on $H_{\bullet}(Z)$. As cohomology is dual to homology, the refined intersection form may be represented by a (symmetric) element of $H^{\bullet}(Z)\otimes H^{\bullet}(Z)$.
It is useful to relate this concept to the long exact sequence in $\mathrm{BM}$-homology:
$\dots\to H^{BM}_{n}(Z)\longrightarrow H^{BM}_{n}(X)\longrightarrow H^{BM}_{n}(X,Z)\to\dots$
If we assume in addition that $Z$ is a deformation retract of $X$, and thus, $H^{\bullet}(X)=H^{\bullet}(Z)$ and $H_{\bullet}(X)=H_{\bullet}(Z)$, then
  • since $Z$ is compact, we have $H^{BM}_{n}(Z)=H_{n}(Z)$;
  • by Poincaré duality $H^{BM}_{n}(X)\cong H^{2m-n}(X)$ and by our assumption we have $H^{2m-n}(X)\cong H^{2m-n}(Z)$;
  • by Lefschetz duality $H^{BM}_{n}(X,Z)\cong H^{2m-n}(X\setminus Z)$.
We can now rewrite our sequence as the sequence
(69) $\dots\to H_{n}(Z)\overset{i_{*}}{\longrightarrow}H^{2m-n}(Z)\overset{j^{*}}{\longrightarrow}H^{2m-n}(X\setminus Z)\to\dots$
which is exact in the middle, i.e. $\operatorname{im}i_{*}\cong\ker j^{*}$. Now we can conclude the following:
Proposition 7.1.
Let $X$ be a smooth variety, and $Z\subset X$ a projective subvariety, which is a deformation retract of $X$. Denote by $L$ the direct sum of the kernels of the map $j^{*}$ in (69) over all degrees. Then
  1. (1)
    $j^{*}$ is the cohomological restriction map, and, in particular, $L\subset H^{\bullet}(Z)$ is an ideal.
  2. (2)
    The map $i_{*}$ in (69) is induced by the refined intersection form on $H_{n}(Z)\otimes H_{2m-n}(Z)$.
  3. (3)
    One can characterize $L^{2m-n}$, the degree-(2m-n) part of $L$, as the annihilator of the radical of the refined intersection form: $w\in L^{2m-n}$ if and only if $w\cap T=0$ for all $T\in H_{2m-n}(Z)$ such that $[S]\cdot_{X}[T]=0$ for all $S\in H_{n}(Z)$.
Corollary 7.2.
Assume that there are $\alpha,\beta\in H^{\bullet}(Z)$ such that the refined intersection form is $\alpha\otimes\beta\in H^{\bullet}(Z)\otimes H^{\bullet}(Z)$. Then $\alpha,\beta\in\operatorname{im}i_{*}\cong L.$
Indeed, we can choose a homology class $B$ of $Z$, such that $\beta(B)\neq 0$, and, according to Proposition 7.1(2), $i_{*}B$ is a multiple of $\alpha$.
7.2. The refined intersection form for abelian points
Now let $V\in M_{\rho}^{\mathrm{ab}}$ be an abelian point in the stratum $M_{\rho}^{\mathrm{ab}}$ and recall the notation $\mathcal{C}_{\rho}$ for the normal cone of $M_{\rho}^{\mathrm{ab}}$ in $\mathcal{M}_{0}(r)$ at $V$. According to Theorem 1.7, $L_{\rho}$ may be obtained as the kernel of a certain cohomological restriction map (6). We note that for a small neighborhood $U$ of a generic $V\in M_{\rho}^{\mathrm{ab}}$, the pair $(\pi^{-1}(U),\pi^{-1}(U\setminus M_{\rho}^{\mathrm{ab}}))$ retracts to the pair $(\pi^{-1}(\mathcal{C}_{\rho}),\pi^{-1}(\mathcal{C}_{\rho}\setminus\{V\}))$, and thus we can represent $L_{\rho}$ as the kernel of the map
(70) $j^{*}:H^{\bullet}(F_{\rho})\cong H^{\bullet}(\pi^{-1}(\mathcal{C}_{\rho}))\longrightarrow H^{\bullet}(\pi^{-1}(\mathcal{C}_{\rho}\setminus\{V\}))$
between the cohomology rings of two smooth manifolds. Now, according to §7.1, the pair
$X=\pi^{-1}(\mathcal{C}_{\rho})\quad\text{and}\quad Z=F_{\rho}$
has a refined intersection form, which may be represented as an element of $H^{\bullet}(F_{\rho})\otimes H^{\bullet}(F_{\rho})$.
Theorem 7.3.
Consider the element
[…]
where $d(\rho)=\sum_{i<j}r_{i}r_{j}(g-1)$ and $x_{i}\in H^{2}(F_{\rho})$, $i=1,\ldots,k$, are the cohomology classes defined in Lemma 4.15. The refined intersection form on
$H_{2d(\rho)}(F_{\rho})\otimes H_{2(d(\rho)+r-k)}(F_{\rho})\text{ is equal to }(-1)^{d(\rho)}w_{\rho}\otimes w_{\rho}\prod_{i=1}^{k}x_{i}^{r_{i}-1}.$
Before presenting the proof, we need to make a small detour.
7.3. Interlude on excess intersection formulas
Our plan is to calculate all intersections of all homology classes of $F_{\rho}$ in degrees $2d(\rho)$ and $2(d(\rho)+r-k)$, respectively. We note that $H_{2(d(\rho)+r-k)}(F_{\rho})$ has a natural basis: the fundamental classes of the $k!$ irreducible components of $F_{\rho}$.
By the nature of the problem, the calculation of the refined intersection form (68) involves calculating an excess intersection number of the varieties $S$ and $T$ inside $X$. Informally, this means that we need to move $T$ inside $X$ until it becomes transversal to $S$, and then count the intersection points with signs.
We are fortunate, that, even though $F_{\rho}$ is singular, its homology has a basis of smooth cycles, thus our calculations can be reduced to finding excess intersection numbers of smooth subvarieties. These numbers, in turn, are certain Chern numbers of the relevant varieties.
Theorem 7.4.
[13, §13.3] Let $S$ and $T$ be smooth projective subvarieties of a smooth variety $X$ with $\dim S+\dim T=\dim X$. Then
(71) $[S]\cdot_{X}[T]=\int_{S\cap T}\frac{c(N_{S|X})}{c(N_{S\cap T|T})},$
where $c(.)$ is the total Chern class starting with 1.
Assume now that $Z$ is smooth, proper of codimension $m$ in a smooth variety $X$, and $S,T\subset Z$ are smooth, generically transversal with $\dim S+\dim T=\dim X$. Then (71) simplifies to
$[S]\cdot_{X}[T]=\int_{S\cap T}c_{m}(N_{Z|X}).$
In particular, if $T=Z$, then
(72) $[S]\cdot_{X}[Z]=\int_{S}c_{m}(N_{Z|X}).$
7.4. Proof of Theorem 7.3
Proof.
We need to prove an identity for pairs of cycles $S,T$ in $F_{\rho}$ of dimensions $d(\rho)$ and $d(\rho)+r-k$ respectively. Note that necessarily, $T$ will be the fundamental class of one of the projective towers. We separate two cases.
Case 1. $S$ and $T$ belong to the same tower $\mathcal{T}_{\sigma}$. Then we can use (72), where $Z=[\mathcal{T}_{\sigma}]$ and $S$ is a cycle of dimension $d(\rho)$. We introduce the notation $w_{\rho}^{+}=(-1)^{d(\rho)}w_{\rho}\prod_{i=1}^{k}x_{i}^{r_{i}-1}$; to prove Theorem 7.3, we need to show that
$[S]\cdot_{X}[\mathcal{T}_{\sigma}]=\int_{S}w_{\rho}\cdot\int_{\mathcal{T}_{\sigma}}w_{\rho}^{+}.$
According to (72), the left-hand side is equal to $\int_{S}c_{d(\rho)}(N_{\mathcal{T}_{\sigma}|X})$. By Corollary 4.8, the right-hand side is $(-1)^{l(\sigma)}\int_{S}w_{\rho}$, where $l(\sigma)$ is equal to the number of pairs $i<j$, such that $\sigma(i)>\sigma(j)$, counted with multiplicity $r_{\sigma(i)}r_{\sigma(j)}(g-1)$. Thus the result in this case follows from
Lemma 7.5.
The normal bundle of $\mathcal{T}_{\sigma}$ in $X$ has the top Chern class $(-1)^{l(\sigma)}w_{\rho}$.
Proof.
A simple calculation, similar to that of Proposition 4.5, shows that the Chern character of the tangent bundle of the normal slice $X=\pi^{-1}(\mathcal{C}_{\rho})$ is equal to
$ch(TX|_{\mathcal{T}_{\sigma}})=\sum_{i=1}^{k}r_{i}\exp(x_{i})+\sum_{1\leq i\neq j\leq k}r_{i}r_{j}(g-1)\exp(x_{i}-x_{j})-k.$
It follows from Proposition 4.9 that
$ch(T\mathcal{T}_{\sigma})=\sum_{i=1}^{k}r_{i}\exp(x_{i})+\sum_{1\leq i<j\leq k}r_{\sigma(i)}r_{\sigma(j)}(g-1)\exp(x_{\sigma(j)}-x_{\sigma(i)})-k,$
and thus we have
(73) $ch(N_{\mathcal{T}_{\sigma}|X})=(g-1)\sum_{1\leq i<j\leq k}r_{\sigma(i)}r_{\sigma(j)}\exp(x_{\sigma(i)}-x_{\sigma(j)}).$
Hence the top Chern class of the normal bundle of $\mathcal{T}_{\sigma}$ in $X$ is equal to $(-1)^{l(\sigma)}w_{\rho}$. ∎
Case 2. $S$ belongs to the tower $\mathcal{T}_{\sigma}$, while $T$ is $\mathcal{T}_{\tau}$. Then we are in the situation of (71).
Case 2A. Let us first look at the simple case when $\mathcal{T}_{\sigma}$ and $\mathcal{T}_{\tau}$ intersect transversally; in this case the intersection is $\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}=\mathcal{I}_{\square}\cong\prod_{i=1}^{k}\mathbb{P}^{r_{i}-1}$ and for any $1\leq i\neq j\leq k$ we have $\sigma(i)<\sigma(j)\iff\tau(i)>\tau(j)$ (see §4.4). Thus we have the equality
$[S]\cdot_{X}[\mathcal{T}_{\tau}]=[S]\cdot_{\mathcal{T}_{\sigma}}[\mathcal{I}_{\square}].$
Now we need to show that
(74) $[S]\cdot_{\mathcal{T}_{\sigma}}[\mathcal{I}_{\square}]=\int_{S}w_{\rho}\cdot\int_{\mathcal{T}_{\tau}}w_{\rho}^{+}=(-1)^{l(\tau)}\int_{S}w_{\rho}=(-1)^{d(\rho)+l(\sigma)}\int_{S}w_{\rho},$
and this is clearly equivalent to the statement that the Poincaré dual class of $\mathcal{I}_{\square}$ in $\mathcal{T}_{\sigma}$ is $(-1)^{d(\rho)+l(\sigma)}w_{\rho}$, which follows from the integral formula, Corollary 4.8.
Case 2B. The general case is similar: recall the notation $\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}=\mathcal{I}(\sigma,\tau)$. We first evaluate the intersection of these two towers, thinking of this as the element of the homology ring of $\mathcal{I}(\sigma,\tau)$:
$[\mathcal{T}_{\sigma}]\cdot_{X}[\mathcal{T}_{\tau}]=\int_{\mathcal{I}(\sigma,\tau)}w_{\sigma,\tau},$
where $w_{\sigma,\tau}$ is the top Chern class of the bundle $N_{\mathcal{T}_{\sigma}|X}/N_{\mathcal{I}(\sigma,\tau)|\mathcal{T}_{\tau}}.$ Now the statement reduces to the following fact:
Lemma 7.6.
The Poincaré dual of $\mathcal{I}_{\square}$ in $\mathcal{I}(\sigma,\tau)$ is the class $(-1)^{l(\sigma)+l(\tau)+d(\rho)}w_{\sigma,\tau}$.
Proof.
For any $1\leq i\neq j\leq k$ and $\sigma\in\Sigma_{k}$, we will write $i\prec_{\sigma}j$ if $\sigma^{-1}(i)<\sigma^{-1}(j)$.
It follows from Proposition 4.19 that
$ch(T\mathcal{I}(\sigma,\tau))-ch(T\mathcal{I}_{\square})=\sum_{\begin{subarray}{c}i\prec_{\tau}j\\ i\prec_{\sigma}j\end{subarray}}r_{i}r_{j}(g-1)\exp(x_{j}-x_{i}),$
and thus the Poincaré dual element of $\mathcal{I}_{\square}$ in $\mathcal{I}(\sigma,\tau)$ is the class $\prod(x_{j}-x_{i})^{r_{i}r_{j}(g-1)},$ where the product runs over the pairs $i,j$ such that $\tau^{-1}(i)<\tau^{-1}(j)$ and $\sigma^{-1}(i)<\sigma^{-1}(j)$.
Now we will evaluate the element $w_{\sigma,\tau}$. Using Propositions 4.9 and 4.19 we obtain that
$ch(N_{\mathcal{I}(\sigma,\tau)|\mathcal{T}_{\tau}})=\sum_{1\leq i\leq k}r_{i}\exp(x_{i})-k+\sum_{1\leq i<j\leq k}r_{\tau(i)}r_{\tau(j)}(g-1)\exp(x_{\tau(j)}-x_{\tau(i)})-\\ \sum_{1\leq i\leq k}r_{i}\exp(x_{i})+k-\sum_{\begin{subarray}{c}i\prec_{\tau}j\\ i\prec_{\sigma}j\end{subarray}}r_{i}r_{j}(g-1)\exp(x_{j}-x_{i})=\sum_{\begin{subarray}{c}i\prec_{\tau}j\\ j\prec_{\sigma}i\end{subarray}}r_{i}r_{j}(g-1)\exp(x_{j}-x_{i}).$
This together with (73) shows that
$w_{\sigma,\tau}=\prod_{\begin{subarray}{c}j\prec_{\tau}i\\ j\prec_{\sigma}i\end{subarray}}(x_{j}-x_{i})^{r_{i}r_{j}(g-1)}.$
A simple calculation shows that the number of pairs $i,j$ such that $\tau^{-1}(i)>\tau^{-1}(j)$ and $\sigma^{-1}(i)>\sigma^{-1}(j)$, counted with multiplicity $r_{i}r_{j}(g-1)$, has the same parity as $l(\sigma)+l(\tau)+d(\rho)$, which completes the proof of Lemma 7.6. ∎
In other words, we showed that for any two permutations $\sigma$ and $\tau$ in $\Sigma_{k}$, we have
$[\mathcal{T}_{\sigma}]\cdot_{X}[\mathcal{T}_{\tau}]=(-1)^{l(\sigma)+l(\tau)+d(\rho)}[\mathcal{I}_{\square}],$
which implies the equality
$[S]\cdot_{X}[\mathcal{T}_{\tau}]=(-1)^{l(\sigma)+l(\tau)+d(\rho)}[S]\cdot_{\mathcal{T}_{\sigma}}[\mathcal{I}_{\square}].$
Since the Poincaré dual class of $\mathcal{I}_{\square}$ in $\mathcal{T}_{\sigma}$ is $(-1)^{d(\rho)+l(\sigma)}w_{\rho}$ (see Corollary 4.8), we have
$[S]\cdot_{X}[\mathcal{T}_{\tau}]=(-1)^{l(\tau)}\int_{S}w_{\rho}=\int_{S}w_{\rho}\cdot\int_{\mathcal{T}_{\tau}}w_{\rho}^{+}.$
Now we are ready to identify the stalk of the local system $\mathcal{L}_{\rho}$.
Theorem 7.7.
The subspace $L_{\rho}\subset H^{\bullet}(F_{\rho})$ is the ideal generated by the element $w_{\rho}$, and has a canonical basis
[…]
Proof.
It follows from Corollary 7.2 and Theorem 7.3 that $w_{\rho},w_{\rho}\prod_{i=1}^{k}x_{i}^{r_{i}-1}\in L_{\rho}$, and since $L_{\rho}$ is an ideal in $H^{\bullet}(F_{\rho})$, this means that the vector space on the right side of (75) is a subspace of $L_{\rho}$.
To prove the linear independence of the basis elements, we observe that it follows from Corollary 4.8 that for $w_{\rho}\prod_{i=1}^{k}x_{i}^{m_{i}}\in H^{\bullet}(F_{\rho})$ appearing in (75), the monomial $y=\prod_{i=1}^{k}x_{i}^{r_{i}-m_{i}-1}$ has the property
$y\cdot w_{\rho}\prod_{i=1}^{k}x_{i}^{m^{\prime}_{i}}=\begin{cases}w_{\rho}\prod_{i=1}^{k}x_{i}^{r_{1}-1},&\text{\, if \,}m^{\prime}_{i}=m_{i}\;\>\forall i\in\{1,...,k\};\\ 0,&\text{\; if \;}\sum_{i=1}^{k}m^{\prime}_{i}=\sum_{i=1}^{k}m^{\prime}_{i}\text{\; and \;}\exists i:m^{\prime}_{i}\neq m_{i}.\end{cases}$
We thus conclude that the Hilbert function of the vector space on the right side of (75) coincides with the result calculated in Theorem 6.1; hence we arrive at the equality (75). ∎
8. Application of the global Decomposition Theorem, and conclusions
To summarize our results so far: we have shown that for our map $\pi:\mathcal{P}_{0}(r)\to\mathcal{M}_{0}(r)$, we have a decomposition
(76) $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}=\bigoplus\limits_{\rho\vdash r}IC\left(\overline{M}_{\rho},\mathcal{L}_{\rho}\right),$
where $\mathcal{L}_{\rho}$ is the local system whose stalks we identified in Theorem 7.7. In this last section of the paper, we will convert (76) into a recursive statement, which then will lead to our main result (5).
8.1. Quotients and local systems
Our first step is to calculate the monodromy of the local system $\mathcal{L}_{\rho}$. For a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$, we introduce the subgroup $\mathrm{Aut}(\rho)\subset\Sigma_{k}$:
$\mathrm{Aut}(\rho)=\{\sigma\in\Sigma_{k}|\,r_{i}=r_{\sigma(i)},\,i=1,\dots,k\}.$
and the subvariety
$\widetilde{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},V_{2},\dots,V_{k})|\,V_{i}\in\mathcal{M}_{0}(r_{i})^{\mathrm{st}},V_{i}\not\cong V_{j}\text{ if }i\neq j\}\subset\\ \mathcal{M}_{0}(r_{1})^{\mathrm{st}}\times\mathcal{M}_{0}(r_{2})^{\mathrm{st}}\times\dots\times\mathcal{M}_{0}(r_{k})^{\mathrm{st}},$
such that $M_{\rho}^{\mathrm{ab}}=\widetilde{M}_{\rho}^{\mathrm{ab}}/\mathrm{Aut}(\rho)$.
Note that $\mathrm{Aut}(\rho)$ acts on $L_{\rho}$ given in (75) by permuting the variables $x_{i}$, $i=1,2,\dots,k$. Thus we can identify the local system $\mathcal{L}_{\rho}$ over $M_{\rho}^{\mathrm{ab}}$ as
$\mathcal{L}_{\rho}=\widetilde{M}_{\rho}^{\mathrm{ab}}\times_{\mathrm{Aut}(\rho)}L_{\rho}.$
Now we pass to the intermediate extensions. We will need the following basic facts from the theory (see [14, §4.2] and [20, Lemma 2.12]).
Proposition 8.1.
  1. (1)
    Let $\tau:\widetilde{X}\to X$ be the normalization of the projective variety $X$. Then this is an isomorphism on the smooth part $X^{\circ}\subset X$, and we have $\tau_{*}IC(\widetilde{X},\mathbb{Q})=IC(X,\mathbb{Q})$.
  2. (2)
    Assume that a finite group $G$ acts on a variety $X$, and let $R$ be a representation of $G$. Let $X^{\circ}\subset X$ be the part of the smooth locus of $X$ on which $G$ acts freely. Then
    $IH^{\bullet}(X,R)^{G}=IH^{\bullet}(X/G,\mathcal{R}),\text{ where }\mathcal{R}=X^{\circ}\times_{G}R.$
Now introduce the notation
$\mathcal{M}_{\rho}=\mathcal{M}_{0}(r_{1})\times\mathcal{M}_{0}(r_{2})\times\dots\mathcal{M}_{0}(r_{k}).$
Part (2) of Proposition 8.1 implies that
(77) $IH^{\bullet}(\mathcal{M}_{\rho},L_{\rho})^{\mathrm{Aut}(\rho)}=IH^{\bullet}(\mathcal{M}_{\rho}/\mathrm{Aut}(\rho),\mathcal{L}_{\rho}).$
Consider the equivalence relation on $\mathcal{M}_{\rho}$: $(V_{1},\dots,V_{k})\sim(W_{1},\dots,W_{k})$ if
  • $V_{i}\cong W_{i}$ if $V_{i}$ is stable, $i=1,\dots,k$, and
  • $V_{1}\oplus V_{2}\oplus\dots\oplus V_{k}=W_{1}\oplus W_{2}\oplus\dots\oplus W_{k}$.
Example 2.
Let $r_{1},r_{2}=2,$ and $L_{1},L_{2},L_{3},L_{4}$ be 4 different line bundles. Then
$(L_{1}\oplus L_{2},L_{3}\oplus L_{4})\sim(L_{1}\oplus L_{3},L_{2}\oplus L_{4}).$
Denote the quotient of $\mathcal{M}_{\rho}$ with respect to the equivalence relation by $Q_{\rho}=\mathcal{M}_{\rho}/\sim$. Then
  • The canonical quotient map $\mathcal{M}_{\rho}\to Q_{\rho}$ is a normalization.
  • There is an induced action of $\mathrm{Aut}(\rho)$ on $Q_{\rho}$, and $Q_{\rho}/\mathrm{Aut}(\rho)\cong\overline{M}_{\rho}$.
With these preparations, we are ready to formulate the first version of our main result.
Proposition 8.2.
We have the decomposition
(78) $H^{\bullet}(\mathcal{P}_{0}(r))\cong\bigoplus_{\rho\vdash r}\left[IH^{\bullet}(\mathcal{M}_{0}(r_{1}))\otimes IH^{\bullet}(\mathcal{M}_{0}(r_{2}))\otimes\ldots\otimes IH^{\bullet}(\mathcal{M}_{0}(r_{k}))\otimes L_{\rho}\right]^{\mathrm{Aut}(\rho)}.$
Proof.
Indeed, taking global cohomology in (76) (cf. (1.7.C)), we obtain
$H^{\bullet}(\mathcal{P}_{0}(r))\cong\bigoplus_{\rho\vdash r}IH(\overline{M}_{\rho},\mathcal{L}_{\rho}).$
According to Proposition 8.1, we have
$IH^{\bullet}(\overline{M}_{\rho},\mathcal{L}_{\rho})\cong IH^{\bullet}(Q_{\rho},L_{\rho})^{\mathrm{Aut}(\rho)}\cong IH^{\bullet}(\mathcal{M}_{\rho},L_{\rho})^{\mathrm{Aut}(\rho)}.$
Finally, applying the Künneth formula, we arrive at (78). ∎
At this point, it is more convenient to parameterize the partition $\rho$ using multiplicities: we will rewrite $r=r_{1}+\ldots+r_{k}$ as
$r=\sum_{a=1}^{l}k_{a}n_{a},\quad n_{1}>n_{2}>\ldots>n_{l},$
where $n_{a}$ is one of the elements of the partition $\rho$, and $k_{a}$ is the number of times this element appears in $\rho$. We can then partition the index set $\{1,2,\ldots,k\}$ into $l$ groups $I_{1},\ldots,I_{l}$, such that $|I_{a}|=k_{a}$, and $j\in I_{a}$ exactly when $r_{j}=n_{a}$. Then
[…]
is a product of permutation groups, and we can write $w_{\rho}$ (cf. Theorem 7.3) as
(79) $w_{\rho}=\prod_{a<b}\prod_{i\in I_{a},\,j\in I_{b}}(x_{i}-x_{j})^{n_{a}n_{b}(g-1)}\times\\ \prod_{i,j\in I_{1},\,i<j}(x_{i}-x_{j})^{n_{1}^{2}(g-1)}\times\ldots\times\prod_{i,j\in I_{l},\,i<j}(x_{i}-x_{j})^{n_{l}^{2}(g-1)}.$
We note that the subspace $\mathbb{Q}w_{\rho}\in L_{\rho}$ is invariant under the $\mathrm{Aut}(\rho)$-action and our immediate goal is to identify this 1-dimensional representation in the category of graded representations.
[…]
$\mathbb{S}^{n}[-q]\cong\begin{cases}\mathbb{Q}[-nq]\text{ if }n\text{ is even},\\ \mathbb{S}[-nq]\text{ if }n\text{ is odd}.\end{cases}$
The first factor of (79) is invariant under $\mathrm{Aut}(\rho)$, and thus, since the variables $x_{i}$ have degree 2, it contributes $\mathbb{Q}[(1-g)(r^{2}-k_{1}^{2}n_{1}^{2}-\ldots-k_{l}^{2}n_{l}^{2})]$. The factor corresponding to $I_{a}$ in (79) is
$\mathbb{S}_{a}^{n_{a}^{2}(1-g)}[k_{a}(k_{a}-1)n_{a}^{2}(1-g)]=\mathbb{S}_{a}^{n_{a}^{2}(1-g)}[k_{a}^{2}n_{a}^{2}(1-g)]\otimes\mathbb{Q}[-k_{a}n_{a}^{2}(1-g)].$
Combining these two identifications, we obtain
Lemma 8.3.
As a 1-dimensional representation of $\mathrm{Aut}(\rho)$,
$\mathbb{Q}w_{\rho}\cong\mathbb{Q}[(1-g)r^{2}]\otimes\mathbb{S}_{1}^{(1-g)n_{1}^{2}}[(g-1)k_{1}n_{1}^{2}]\boxtimes\ldots\boxtimes\mathbb{S}_{l}^{(1-g)n_{l}^{2}}[(g-1)k_{l}n_{l}^{2}].$
Finally, inserting the second factor of (75) in the form $\oplus_{j=0}^{n}\mathbb{Q}x^{j}\cong H^{\bullet}(\mathbb{P}^{n})$, we obtain that, as a graded representation of $\mathrm{Aut}(\rho)$, we have
(80) $L_{\rho}[(g-1)r^{2}]\cong\bigotimes_{a=1}^{l}\mathbb{S}^{(1-g)n_{a}^{2}}_{a}\otimes\left(H^{\bullet}(\mathbb{P}^{n_{a}-1})[(g-1)n_{a}^{2}])\right)^{\otimes k_{a}},$
where $\Sigma_{k_{a}}$ acts on $W^{\otimes k_{a}}$ by permuting the tensor factors.
Now, up to a degree shift of $(g-1)r^{2}$, we can rewrite the right-hand side of (78) as
(81) $\bigotimes_{a=1}^{l}\left[\mathbb{S}^{(1-g)n_{a}^{2}}_{a}\otimes\left(IH^{\bullet}(\mathcal{M}_{0}(n_{a}))\otimes H^{\bullet}(\mathbb{P}^{n_{a}-1})[(g-1)n_{a}^{2}]\right)^{\otimes k_{a}}\right]^{\Sigma_{k_{a}}}.$
Remark 8.4.
Note that there is a subtlety in this last equation: $IH^{\bullet}(\mathcal{M}_{0}(n_{a}))$ has odd elements, and the action of a transposition on $\alpha_{1}\otimes\alpha_{2}$ for odd classes $\alpha_{1},\alpha_{2}$ has a minus sign, according to the usual sign conventions. The shifts in the grading have no effect on parity at this point. In what follows, however, we will consider the polynomial rings over graded vector spaces, and there (cf. (4)) the odd degree elements always anticommute.
Lemma 8.5.
The tensor factor with index $a$ in (81),
[…]
is the ”polynomial degree-$k_{a}$” homogeneous piece in the polynomial ring over the graded vector space
$IH^{\bullet}(\mathcal{M}_{0}(n_{a}))\otimes H^{\bullet}(\mathbb{P}^{n_{a}-1})\left[(g-1)n_{a}^{2}\right].$
Proof.
Indeed, the shift by $(g-1)n_{a}^{2}$ will change parity of the degrees when $g$ is even and $n_{a}$ is odd, and this is reflected in the factor $\mathbb{S}^{(1-g)n_{a}^{2}}_{a}$. ∎
8.2. The plethystic exponential and the final result
Now summing up (81) over all partitions $\rho$ of the integer $r$, and taking into account the shift on the left hand side of (80), we arrive at the following interpretation of the right hand side of (78).
Corollary 8.6.
Consider the bigraded vector space
$\mathcal{H}_{0}=q\cdot IH^{\bullet}(\mathcal{M}_{0}(1))\left[g-1\right]\oplus\ldots\oplus q^{r}\cdot IH^{\bullet}(\mathcal{M}_{0}(r))\otimes H^{\bullet}(\mathbb{P}^{r-1})\left[(g-1)r^{2}\right]\oplus\ldots,$
where $q^{j}\cdot W$ is a mnemonic for setting the $q$-degree of $W$ to $j$; the usual cohomological degree will be called the $t$-degree. Then the bigraded vector space
(82) $\bigoplus_{\rho\vdash r}\left[IH^{\bullet}(\mathcal{M}_{0}(r_{1}))\otimes IH^{\bullet}(\mathcal{M}_{0}(r_{2}))\otimes\ldots\otimes IH^{\bullet}(\mathcal{M}_{0}(r_{k}))\otimes L_{\rho}[(g-1)r^{2}]\right]^{\mathrm{Aut}(\rho)}$
is isomorphic to the $q$-degree=$r$ piece of the polynomial ring $\mathbb{Q}_{[t]}[\mathcal{H}_{0}]$ on the vector space $\mathcal{H}_{0}$ defined in (4).
Finally, we sum the left hand side of (78) shifted by $[(g-1)r^{2}]$ over all values of the rank $r$: we consider the bigraded vector space
$\mathcal{H}_{1}=\mathbb{C}\oplus q\cdot H^{\bullet}(\mathcal{P}_{0}(1))[g-1]\oplus\ldots\oplus q^{r}\cdot H^{\bullet}(\mathcal{P}_{0}(r))[(g-1)r^{2}]\oplus\ldots.$
Recall from Corollary 3.12 and Proposition 3.10 that $\mathcal{P}_{0}(r)\cong\mathcal{P}_{1}(r)$ and that $\mathcal{P}_{1}(r)\to\mathcal{M}_{1}(r)$ is a projective bundle with fibers $\mathbb{P}^{r-1}$, hence we have $P_{t}(\mathcal{P}_{0}(r))=P_{t}(\mathcal{M}_{1}(r))\cdot P_{t}(\mathbb{P}^{r-1})$, and thus we can write
$\mathrm{Hilb}_{t,q}(\mathcal{H}_{1})=1+\sum_{r=1}^{\infty}P_{t}(\mathbb{P}^{r-1})\cdot P_{t}(\mathcal{M}_{1}(r))\cdot t^{(1-g)r^{2}}q^{r}.$
Next, combining Proposition 8.2 and Corollary 8.6, we arrive at our main result:
Theorem 8.7.
In the notation introduced in §1.4.2, we have
(83) $\mathrm{Hilb}_{-t,q}(\mathcal{H}_{1})=\mathrm{Hilb}_{-t,q}(\mathbb{Q}_{[t]}[\mathcal{H}_{0}])=\mathbb{E}\mathrm{xp}\left[\mathrm{Hilb}_{-t,q}(\mathcal{H}_{0})\right],$
More explicitly,
(84) $1+\sum_{r=1}^{\infty}P_{-t}(\mathbb{P}^{r-1})\cdot P_{-t}(\mathcal{M}_{1}(r))\cdot(-t)^{(1-g)r^{2}}q^{r}=\\ =\mathbb{E}\mathrm{xp}\left[\sum_{r=1}^{\infty}P_{-t}(\mathbb{P}^{r-1})\cdot IP_{-t}(\mathcal{M}_{0}(r))\cdot(-t)^{(1-g)r^{2}}q^{r}\right].$
Remark 8.8.
Clearly, formula (84), provides us with a recursion for $IP_{-t}(\mathcal{M}_{0}(r))$, if $P_{-t}(\mathcal{M}_{1}(r))$ is known for all $r$.
8.3. Variants and refinements
Now we rewrite this recursion in a form that is even more symmetric, but hides the shifts/parity changes to some extent.
For a compact complex manifold of complex dimension $n$, we can introduce a symmetrized version of the Poincaré polynomial:
$\widehat{P}_{t}(M)=t^{-n}\cdot P_{t}(M)\quad\;\text{satisfying}\quad\;{\widehat{P}}_{1/t}(M)={\widehat{P}}_{t}(M);$
similarly, we introduce the notation $I\widehat{P}_{t}(M)=t^{-n}\cdot IP_{t}(M)$ for the case when $M$ is singular.
Using the dimension formula $\dim\mathcal{M}_{0}(r)=\dim\mathcal{M}_{1}(r)=(g-1)r^{2}+1$ and substituting $\hat{q}=tq$, we can rewrite (84) in the following elegant form:
$1+\sum_{r=1}^{\infty}\widehat{P}_{-t}(\mathbb{P}^{r-1})\cdot\widehat{P}_{-t}(\mathcal{M}_{1}(r))\cdot(-\hat{q})^{r}=\\ =\mathbb{E}\mathrm{xp}\left[\sum_{r=1}^{\infty}\widehat{P}_{-t}(\mathbb{P}^{r-1})\cdot I\widehat{P}_{-t}(\mathcal{M}_{0}(r))\cdot(-\hat{q})^{r}\right].$
Finally, we show that our formulas can be generalized to describe the Hodge structuure on $IH^{\bullet}(\mathcal{M}_{0}(r))$. For a smooth, projective variety $Y$, denote the generator of its Hodge numbers by
$h_{u,v}(Y)=\sum_{p,q}(-1)^{p+q}\dim(H^{p,q}(Y))u^{p}v^{q}.$
Because of purity, setting $u=v=t$, we recover the E-polynomial (cf. 3.18): $E_{t}(Y)=h_{t,t}(Y)$
Similarly, for $Y$ projective, possibly singular, the intersection cohomology is pure, and we can thus introduce the analogous generating function:
$Ih_{u,v}(Y)=\sum_{p,q}(-1)^{p+q}\dim(IH^{p,q}(Y))u^{p}v^{q}.$
Since our arguments were compatible with Hodge structures, we can refine (84) to the equality
(85) $1+\sum_{r=1}^{\infty}h_{u,v}(\mathbb{P}^{r-1})\cdot h_{u,v}(\mathcal{M}_{1}(r))\cdot(-uv)^{(1-g)r^{2}}q^{r}=\\ =\mathbb{E}\mathrm{xp}\left[\sum_{r=1}^{\infty}h_{u,v}(\mathbb{P}^{r-1})\cdot Ih_{u,v}(\mathcal{M}_{0}(r))\cdot(-uv)^{(1-g)r^{2}}q^{r}\right].$